### v2

# On the maximal correlation of some stochastic processes

Abstract
We consider the maximal correlation coefficient $R(X,Y)$ between two stochastic processes $X$ and $Y$. When $(X,Y)$ is a random walk, the result is a consequence of Csáki-Fischer identity and lower-semi continuity of $\text{Law}(X,Y)\to R(X,Y)$. When $(X,Y)$ are two-dimensional Lévy processes, we give an expression of $R(X,Y)$ via the covariance $\Sigma$ and the Lévy measure $\nu$ appeared in the Lévy-Khinchine formula. As a consequence, for two-dimensional $\alpha$-stable random variables $(X,Y)$ with $20<\alpha<$ 2, we give an expression of $R(X,Y)$ via the stability index $\alpha$ and the spectral measure $\tau$ of the $\alpha$-stable distribution. We also prove analogs and generalizations of Dembo-Kagan-Shepp-Yu inequality and Madiman-Barron inequality. Roughly speaking, we investigate the maximal correlation coefficient between two random selected “subvectors” $Y$ and $Z$ of a common random vector $X$. Besides, by using above new results, we recover several classical results.
1 Introduction
For two non-degenerate real random variables $X$ and $Y$, the Pearson correlation coefficient $\rho(X,Y)$ is defined by
$\rho(X,Y)=\frac{\operatorname{Cov}(X,Y)}{\sqrt{\operatorname{Var}(X)}\sqrt{\operatorname{Var}(Y)}}.$
It measures the linear dependence of $X$ and $Y$. It is well-known that $\rho(X,Y)=0$ if $X$ and $Y$ are independent. However, the converse is not true in general. For two non-degenerate random variables $X$ and $Y$, the maximal correlation coefficient between $X$ and $Y$, introduced by Gebelein [Geb41], is given by
$R(X,Y)=\sup\rho(\varphi(X),\psi(Y)),$ (1)
where the supremum is taken for all measurable functions $\varphi$ and $\psi$ such that
$0<E[(\varphi(X))^{2}]<\infty,0<E[(\psi(Y))^{2}]<\infty.$
If $X$ or $Y$ is degenerate, we define $R(X,Y)=0$. The quantity $R(X,Y)$ measures the dependence between $X$ and $Y$. Note that $R(X,Y)=0$ if and only if $X$ and $Y$ are independent, see e.g. [Sar58].
It is well known that the maximal correlation coefficient has the following alternative definition: For a random variable $X$, denote by $L^{2}_{0}(X)$ the Hilbert space of real random variables $Z$ measurable with respect to the sigma-field $\sigma(X)$ generated by $X$ such that $E(Z)=0$ and $E(Z^{2})<\infty$. Then, $R(X,Y)$ equals the operator norm of the conditional expectation $\varphi(X)\mapsto E(\varphi(X)|Y)$ from $L^{2}_{0}(X)$ to $L^{2}_{0}(Y)$. Equivalently, we have that
$R^{2}(X,Y)=\sup\{E[(E(\varphi(X)|Y))^{2}]:E[\varphi(X)]=0,E[\varphi^{2}(X)]=1\}.$ (2)
Recently, the maximal correlation coefficient is generalized and studied by Dadoun and Youssef [DY21] in the context of free probability. But we will concentrate on classical notion of maximal correlation in the present paper.
The maximal correlation coefficient plays an important role in various areas of probability and statistics, including information theory [Cou16, MB07], hyper-contractivity ribbon of a pair of random variables and impossibility of non-interactive simulations of joint distributions [AG76, KA12, KA16], optimal transformation for regression [BF85], the spectral gap of Markov chains and the convergence theory of data augmentation algorithms [LWK94].
As pointed out by Rényi in [R5́9], it is often difficult to calculate the maximal correlation coefficient although it may seem to be not so difficult at the first sight. One reason is that the supremum in (1) could not be replaced by maximum in general. The explicit formula of the maximal correlation coefficient $R(X,Y)$ is known in a few cases. For example, when $(X,Y)$ is jointly Gaussian,
$R(X,Y)=|\rho(X,Y)|,$ (3)
see [Lan57]. If $(X,Y)$ is uniformly distributed in the domain $|x|^{p}+|y|^{q}\leq 1$ with $p,q>0$, then
$R(X,Y)=\frac{1}{\sqrt{(p+1)(q+1)}},$
see [CF60, Example 5 in Section 2]. They also found the maximal correlation for multinominal distributions and multihypergeometric distributions in [CF63, Section 6]. Bücher and Staud [BS25] have obtained the maximal correlation coefficient for bivariate Marshall Olkin exponential distribution. There are also several results for maximal correlation coefficients by using orthogonal polynomials, see [LBCnM06, LBSMn14, LBSMn98, Nev92, PX13, SM85, Ter83]. Among these results, [SM85] gives the maximal correlation coefficients for Dirichlet distributions, which provides sharp upper bounds for the maximal correlation coefficients for order statistics studied in [SM85, Ter83]. In [LBSMn98, Nev92], they studied the maximal correlation between the $i$-th and $j$-th records. In [PX13], they provided a unified approach for obtaining the maximal correlation coefficient for a subclass of Lancaster distributions. For maximal correlation coefficients for general distributions, we are aware of the following results. Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are independent. Then, we have that
$R((X_{1},X_{2}),(Y_{1},Y_{2}))=\max(R(X_{1},Y_{1}),R(X_{2},Y_{2})).$
This is known as Csáki-Fischer identity [CF63, Theorem 6.2], see also [Wit75, Theorem 1]. If $X_{1},X_{2},\ldots,X_{n}$ are independent and identically distributed (i.i.d.) non-degenerate random variables, then for $m\leq n$, we have that
$R(S_{n},S_{m})=\sqrt{m/n},$ (4)
where $S_{k}=\sum_{i=1}^{k}X_{i}$ is the partial sum of $(X_{i})_{i}$, see [BDK05, DKS01, Nov04]. Let’s call the upper bound $R(S_{n},S_{m})\leq\sqrt{m/n}$ the Dembo-Kagan-Shepp (DKS) inequality. Yu [Yu08, Theorem 4.1] further generalized (4) to
$R\left(\sum_{i=1}^{m}X_{i},\sum_{j=\ell+1}^{n}X_{j}\right)=\frac{m-\ell}{\sqrt{m(n-\ell)}}$ (5)
for $1\leq\ell+1\leq m\leq n$.
We are interested in the maximal correlation coefficient between two stochastic processes, e.g. two random walks, two Lévy processes or two random sub-vectors of a third common random vector.
To be more precise, let $(\xi_{n},\eta_{n})_{n\geq 1}$ be i.i.d. random vectors. Let $(S_{n},T_{n})_{n\geq 0}$ be a random walk starting from 00 with increments $(\xi_{n},\eta_{n})$, i.e.
$(S_{0},T_{0})=0\text{ and for }n\geq 1,(S_{n},T_{n})=\sum_{m=1}^{n}(\xi_{m},\eta_{m}).$ (6)
Let $S=(S_{n})_{n\geq 0}$ and $T=(T_{n})_{n\geq 0}$. Then, we have the following result:
Theorem 1.1 (Csáki-Fischer).
$R((S_{n})_{n\leq m},(T_{n})_{n\leq m})=R(\xi_{1},\eta_{1})=R(S,T).$
The first equality is a direct consequence of Csáki-Fischer identity (Theorem 2.2) since
$R((S_{n})_{n\leq m},(T_{n})_{n\leq m})=R((\xi_{1},\xi_{2},\ldots,\xi_{m}),(\eta_{1},\eta_{2},\ldots,\eta_{m})).$
For the second equality, we use the lower semi-continuity of $\text{Law}(X,Y)\mapsto R(X,Y)$ (Lemma 2.3). For self-containedness, we also provide a new probabilistic proof of the Csáki-Fischer identity. As a consequence of Theorem 1.1, the central limit theorem and the lower semi-continuity of $\text{Law}(X,Y)\mapsto R(X,Y)$ (Lemma 2.3), we give a new proof of (3). Similarly, by using the law of rare events instead of the central limit theorem, we see that (3) holds for bivariate Poisson distributions.
Next, we move to a new related problem. We have considered the maximal correlation coefficient for random walks. As Lévy processes are continuous analogs of random walks, it is natural to consider the maximal correlation coefficient between two correlated Lévy processes $(X_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$. We assume that $(X_{t},Y_{t})$ is a two dimensional Lévy process. We refer to Applebaum’s book [App09] for properties of Lévy processes. The important Lévy-Khinchine formula is well-known. In the two-dimensional case, it has the following form:
$E[e^{i(u_{1}X_{t}+u_{2}Y_{t})}$ $]=\exp\left(t\left\{i(b_{1}u_{1}+b_{2}u_{2})-\frac{1}{2}\sum_{j,k=1,2}\Sigma_{jk}u_{j}u_{k}\right.\right.$
$+\left.\left.\int_{\mathbb{R}^{2}\setminus\{0\}}\left(e^{i(u_{1}x+u_{2}y)}-1-i(u_{1}x+u_{2}y)I_{(0,1]}(x^{2}+y^{2})\right)\,\nu(\mathrm{d}x,\mathrm{d}y)\right\}\right).$ (7)
Here, $\Sigma=(\Sigma_{jk})_{j,k=1,2}$ is the covariance matrix of the Brownian motion in the Lévy-Itô decomposition, and $\nu(\mathrm{d}x,\mathrm{d}y)$ is the Lévy measure on $\mathbb{R}^{2}\setminus\{0\}$ satisfying
$\int_{\mathbb{R}^{2}\setminus\{0\}}\min(1,x^{2}+y^{2})\,\nu(\mathrm{d}x,\mathrm{d}y)<\infty.$
Theorem 1.2.
Let $(X_{t},Y_{t})_{t\geq 0}$ be a two dimensional Lévy process with the Lévy-Khinchine formula (1). Define
$\rho=\left\{\begin{array}[]{ll}\frac{\Sigma_{12}}{\sqrt{\Sigma_{11}\Sigma_{22}}},&\Sigma_{11}\Sigma_{22}>0,\\ 0,&\Sigma_{11}\Sigma_{22}=0.\end{array}\right.$
Let $\mathrm{Op}(\nu)$ be the minimal constant $s$ such that the inequality
$\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)\leq s\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}$
holds for all measurable functions $\varphi(x)$ and $\psi(y)$ such that $\varphi(0)=\psi(0)=0$. Then, we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})=\max(|\rho|,\mathrm{Op}(\nu)).$
Next, we consider applications of Theorem 1.2 to stable random vectors. Lancaster have obtained an expression of the maximal correlation coefficients for jointly Gaussian random variables, see (3). It is natural to ask for an expression for jointly stable random variables. The index $\alpha$ of stability takes values in $(0,2]$. When $\alpha=2$, it must be jointly Gaussian. Therefore, we consider the case $\alpha\in(0,2)$. Suppose $(X,Y)$ is a two-dimensional stable random vector. Then, it is infinitely divisible and has the following Lévy-Khinchine formula:
$E[e^{i(u_{1}X+u_{2}Y)}$ $]=\exp\Big{(}\Big{\{}i(b_{1}u_{1}+b_{2}u_{2})$
$+\left.\left.\int_{\mathbb{R}^{2}\setminus\{0\}}\left(e^{i(u_{1}x+u_{2}y)}-1-i(u_{1}x+u_{2}y)I_{(0,1]}(x^{2}+y^{2})\right)\,\nu(\mathrm{d}x,\mathrm{d}y)\right\}\right).$ (8)
Here, the Lévy measure $\nu$ takes the form
$\nu(B)=\int_{0}^{\infty}\int_{\mathbb{R}/(2\pi\mathbb{Z})}I_{B}(r\cos\theta,r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta),\quad B\in\mathcal{B}(\mathbb{R}^{2}\setminus\{0\}),$ (9)
where the spectral measure $\tau(\mathrm{d}\theta)$ is a finite Borel measure.
Theorem 1.3.
Let $(X,Y)\in\mathbb{R}^{2}$ be a stable random vector with the index $\alpha\in(0,2)$. Let
$C_{++}=\int_{0}^{\frac{\pi}{2}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),\quad$ $C_{+-}=\int_{\frac{3\pi}{2}}^{2\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),$
$C_{-+}=\int_{\frac{\pi}{2}}^{\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),\quad$ $C_{--}=\int_{\pi}^{\frac{3\pi}{2}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),$
and
$D^{x}_{+}=\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}|\cos\theta|^{\alpha}\,\tau(\mathrm{d}\theta),\quad$ $D^{x}_{-}=\int_{\frac{\pi}{2}}^{\frac{3\pi}{2}}|\cos\theta|^{\alpha}\,\tau(\mathrm{d}\theta),$
$D^{y}_{+}=\int_{0}^{\pi}|\sin\theta|^{\alpha}\,\tau(\mathrm{d}\theta),\quad$ $D^{y}_{-}=\int_{\pi}^{2\pi}|\sin\theta|^{\alpha}\,\tau(\mathrm{d}\theta).$
Then, we have that
$R(X,Y)=\mathrm{Op}(\nu)=\left\|\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}\right\|_{2}$
with the convention that $/0=0$ 0, where $\|\cdot\|_{2}$ means the spectral norm.
For random vectors $(\widetilde{X},\widetilde{Y})$ in the domain of contraction of stable laws, Theorem 1.3 provides a lower bound for $R(\widetilde{X},\widetilde{Y})$, see Remark 4.1 below.
Finally, we consider analogs and generalizations of Dembo-Kagan-Shepp inequality. We start by the following proposition, which is proved by using Csáki-Fischer identity.
Proposition 1.4.
Let $(X_{n})_{n\geq 1}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $(B_{n})_{n\geq 1}$ be i.i.d. Bernoulli random variables such that $P(B_{n}=1)=1-P(B_{n}=0)=p_{n}$ and $(B_{n})_{n\geq 1}$ is independent of $(X_{n})_{n\geq 1}$. Define
$Y_{n}=\left\{\begin{array}[]{ll}X_{n},&\text{if }B_{n}=1,\\ \partial,&\text{if }B_{n}=0.\end{array}\right.$
Then, for $n\geq 1$, we have that
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))=\sqrt{\max(p_{1},p_{2},\ldots,p_{n})}.$
Let $T=\{i=1,2,\ldots,n:B_{n}=1\}$. Then, $T$ is a random subset of $[n]:=\{1,2,\ldots,n\}$. Note that $\max(p_{1},p_{2},\ldots,p_{n})=\max\{i\in[n]:P(i\in T)\}$. Hence, the following theorem due to Madiman and Barron is a generalization of Proposition 1.4.
Theorem 1.5 (Madiman-Barron).
Let $X_{1},X_{2},\ldots,X_{n}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $T$ be a random subset of $[n]=\{1,2,\ldots,n\}$. Suppose that $T$ is independent from $(X_{1},X_{2},\ldots,X_{n})$. For $i=1,2,\ldots,n$, define
$Y_{i}=\left\{\begin{array}[]{ll}X_{i},&\text{if }i\in T,\\ \partial,&\text{otherwise}.\end{array}\right.$
Then, for $n\geq 1$, we have that
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))=\sqrt{\max\{i\in[n]:P(i\in T)\}}.$
The lower bound $R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\geq\sqrt{\max\{i\in[n]:P(i\in T)\}}$ is a simple consequence of the fact $R(X_{i},Y_{i})=\sqrt{P(i\in T)}$, which follows from Proposition 1.4 for $n=1$. The most interesting part is the upper bound
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{\max\{i\in[n]:P(i\in T)\}}.$
This upper bound is known in a different form by Madiman and Barron, see [MB07, Lemmas 2 and 4]. In the same paper, this upper bound is further used to deduce generalized entropy power inequalities, monotonicity of Fisher information and the Fisher information inequality of Stam. It is important follow-up work of the break-through work [ABBN04] by Artstein, Ball, Barthe and Naor on the monotonicity of Shannon entropy.
Remark 1.1.
The random variables $X_{1},X_{2},\ldots,X_{n}$ are not necessarily identically distributed.
Remark 1.2.
The Madiman-Barron inequality is a generalization of Dembo-Kagan-Shepp inequality. Indeed, let $X_{1},X_{2},\ldots,X_{n}$ be independent random variables taking values in a general measurable space $(F,\mathcal{F})$. For fixed $m\in[n]=\{1,2,\ldots,n\}$, let $T$ be a uniform subset of $\{1,2,\ldots,n\}$ with size $m$. For $i=1,2,\ldots,n$, define
$Y_{i}=\left\{\begin{array}[]{ll}X_{i},&\text{if }i\in T,\\ \partial,&\text{otherwise},\end{array}\right.$
where $\partial$ is some special point in $(F,\mathcal{F})$. Then, by the Madiman-Barron inequality, for $n\geq 1$,
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{m/n}.$
Suppose $F$ is a vector space. Then, $(\sum_{i=1}^{n}X_{n},\sum_{i=1}^{n}Y_{n}1_{Y_{n}\neq\partial})$ has the same joint distribution as $(S_{n},S_{m})$ in (4). Hence, we have that
$R(S_{n},S_{m})\leq R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))\leq\sqrt{m/n}.$
Note that Theorem 1.5 gives the maximal correlation coefficient between a finite independent sequence and its random sub-sequence. Hence, it is natural to ask for a generalization for two random sub-sequences of the same finite independent sequence. We provide an answer by the following theorem:
Theorem 1.6.
Let $X_{1},X_{2},\ldots,X_{n}$ be non-degenerate independent random variables taking values in a general measurable space $(F,\mathcal{F})$. Fix a special point $\partial$ outside of $F$. Let $S$ and $T$ be two random subsets of $[n]=\{1,2,\ldots,n\}$. Suppose that $(S,T)$ is independent from $(X_{1},X_{2},\ldots,X_{n})$. For $i=1,2,\ldots,n$, define
$Y_{i}=\left\{\begin{array}[]{ll}X_{i},&\text{if }i\in S,\\ \partial,&\text{otherwise},\end{array}\right.\text{ and }Z_{i}=\left\{\begin{array}[]{ll}X_{i},&\text{if }i\in T,\\ \partial,&\text{otherwise}.\end{array}\right.$
Then, for $n\geq 1$, we have that
$R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))=\max(R(S,T),\max(r_{j}:j\in[n])),$ (10)
where $r_{j}$ is the best constant $r$ for the following inequality
$\sum_{s,t:j\in s\cap t}P(S=s,T=t)\alpha_{s}\beta_{t}\leq r\sqrt{\sum_{s:j\in s}P(S=s)\alpha_{s}^{2}}\sqrt{\sum_{t:j\in t}P(T=t)\beta_{t}^{2}}$ (11)
for arbitrary real constants $\alpha_{s}$ and $\beta_{t}$. In particular, when $S$ and $T$ are independent, we have $R(S,T)=0$ and $r_{j}=\sqrt{P(j\in S)P(j\in T)}$.
Here, the non-degeneracy of $X_{i}$ is crucial for the lower bound
$R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))\geq\max(r_{j}:j\in[n]).$
According to Theorem 1.6, in order to calculate $R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))$, we need to know $R(S,T)$. When $S$ and $T$ are dependent, we have no answer in general. However, we have a partial answer in the following special case.
Theorem 1.7.
Let $T$ be a uniform subset of $[n]=\{1,2,\ldots,n\}$ of size $m$. Given $T$, the random set $S$ is a uniform subset of $T$ of size $k\leq m$. Then, we have that
$R(S,T)=\sqrt{\frac{k(n-m)}{m(n-k)}},$
with the convention that $/0=0$ 0.
Recall that Yu [Yu08, Theorem 4.1] proved (5). We give a generalization of the upper bound $R\left(\sum_{i=1}^{m}X_{i},\sum_{j=\ell+1}^{n}X_{j}\right)\leq\frac{m-\ell}{\sqrt{m(n-\ell)}}$ in (5) as an application of Theorem 1.6:
Corollary 1.8.
We use the notation and assumptions in Theorem 1.6. Assume that $(S,T)$ is uniformly distributed in the following set $\{(s,t):s,t\subset[n],|s|=k,|t|=m,|s\cap t|=\ell\}$. Then, we have that
$R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))=\frac{\ell}{\sqrt{mk}}.$
By adapting the arguments in [Cou16] and [MB07], we find that Theorem 1.6 implies the following results in information theory:
Theorem 1.9.
Consider independent continuous random variables $X_{1},X_{2},\ldots,X_{n}$. Let $S\subset T$ be two nested non-empty random subsets of $[n]=\{1,2,\ldots,n\}$, which are independent from $X=(X_{1},X_{2},\ldots,X_{n})$. Let $R$ be the maximal correlation coefficient between $(S,X_{S})$ and $(T,X_{T})$ given by (10), where $X_{S}=(X_{i})_{i\in S}$ and $X_{T}=(X_{j})_{j\in T}$. Denote by $I(Z)=\int_{-\infty}^{\infty}\frac{(f^{\prime}(z))^{2}}{f(z)}\,\mathrm{d}z$ the Fisher information of the continuous random variable $Z$ with the density $f(z)$. Then, we have that
$\sum_{t\subset[n]}P(T=t)I(\sum_{j\in t}X_{j})\mu_{t}^{2}\leq R^{2}\sum_{s\subset[n]}P(S=s)I(\sum_{i\in s}X_{i})\lambda_{s}^{2},$ (12)
where $\mu_{t}=\sum_{s\subset[n]}P(S=s|T=t)\lambda_{s}$ for $t\subset[n]$ and $(\lambda_{s})_{s\subset[n]}$ is an arbitrary real vector.
Organization of the paper: In Section 2, we present several useful properties of the maximal correlation coefficients, which are used in the following proofs. In Section 3, we consider maximal correlation coefficients for random walks. We prove Theorem 1.1 and give new proof of Lancaster’s classical result (3) in separate subsections. In Section 4, we consider the maximal correlation coefficients for two dimensional Lévy processes, two dimensional stable processes and two dimensional stable distributions. We first give the proof of Theorem 1.2. Then, we use Theorem 1.2 to prove Theorem 1.3. As applications, we also discuss several examples. These are presented in different subsections. In Section 5, we consider Madiman-Barron type equalities and prove Proposition 1.4, Theorem 1.6, Theorem 1.7, Corollary 1.8 and Theorem 1.9 in separate subsections. Among these results, Proposition 1.4 connects Section 3 and Section 5. More precisely, it can be viewed as an application of Theorem 1.1 and it is a special case of Theorem 1.5. The result of the special case in Proposition 1.4 is actually the motivation for our alternative expression for Madiman-Barron equality. (The original description of Madiman and Barron is a bit different.) We also discover a Dembo-Kagan-Shepp-Yu-type upper bound in Proposition 5.5. There, we replace the summation in (5) by minimum in (55). It implies the upper bound in [BS25, Corollary 2.2] for bivariate Marshall Olkin exponential distribution. Finally, we present two open problems in Section 6.
2 Preliminaries
2.1 Basic properties
We collect several basic properties in the following:
  1. 1.
    $R(X,Y)\in[0,1]$.
  2. 2.
    $R(X,Y)=0$ iff $X$ and $Y$ are independent.
  3. 3.
    $R(X,Y)=R(Y,X)$.
  4. 4.
    In general, the supremum in the definition of the maximal correlation coefficient cannot be replaced by the maximum. For instance, suppose that $M$ and $N$ are independent Poisson random variables of parameter $1$. Let $X=M-N$ and $Y=M-2N$. Then, we have that $R(X,Y)=1$. Indeed, $\lim_{m\to\infty}\rho(1_{X=m},1_{Y=m})=1$. However, for any pair of measurable functions $\varphi$ and $\psi$ such that $E\varphi(X)=E\psi(Y)=0$ and $\operatorname{Var}\varphi(X)=\operatorname{Var}\psi(Y)=1$, we have that $\rho(\varphi(X),\psi(Y))<1$. Rényi [R5́9, Theorem 1, Theorem 2] provided sufficient condition under which the maximal correlation could be attained at some $\varphi(X)$ and $\psi(Y)$.
  5. 5.
    If $U=\varphi(X)$ is a measurable function of $X$ and $V=\psi(Y)$ is a measurable function of $Y$, then $R(U,V)\leq R(X,Y)$.
2.2 Submultiplicative property
Lemma 2.1 (Lemma 2.1 of [Yu08]).
Suppose that $X$ and $Z$ are conditionally independent given $Y$, then
$R(X,Z)\leq R(X,Y)R(Y,Z).$
This is the first part of [Yu08, Lemma 2.1]. For self-containedness, we provide a short proof. The result comes from the fact that the operator norm of the composition of two operators is not greater than the product of the operator norms of these two operators. In the proof, the Markov property of $(X,Y,Z)$ is crucially used.
Proof.
By the conditional independence of $X$ and $Z$ given $Y$, the following diagram commutes:
[…]
where $\pi_{1},\pi_{2},\pi_{3},\pi_{4}$ and $\pi_{5}$ are orthogonal projections and $\iota$ is the injection. Equivalently, $\pi_{1}:W\mapsto E(W|Z)$, $\pi_{2}:W\mapsto E(W|Y)$, $\pi_{3}:W\mapsto E(W|Y,Z)$, $\pi_{4}:W\mapsto E(W|Z)$ and $\pi_{5}:W\mapsto E(W|Z)$ are the operators of conditional expectations. Hence, we have that
$R(X,Z)=\|\pi_{1}\|=\|\pi_{4}\circ\pi_{2}\|\leq\|\pi_{4}\|\|\pi_{2}\|=R(Y,Z)R(X,Y),$
where |||||$\cdot$||| | ⋅ | | is the operator norm. ∎
2.3 Csáki-Fischer identity
The following result is known as Csáki-Fischer identity [CF63, Theorem 6.2], see also [Wit75, Theorem 1].
Theorem 2.2 (Csáki-Fischer identity).
Suppose that $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ are independent. Then, we have that
$R((X_{1},X_{2}),(Y_{1},Y_{2}))=\max(R(X_{1},Y_{1}),R(X_{2},Y_{2})).$
We provide a simple probabilistic proof for self-containedness.
Proof.
Let $\mathcal{F}=\sigma(X_{1},Y_{1})$. Note that
$\mathrm{Cov}(f(X_{1},X_{2}),g(Y_{1},Y_{2}))$ $=E[\mathrm{Cov}(f(X_{1},X_{2}),g(Y_{1},Y_{2})|\mathcal{F})]$
$\quad+\mathrm{Cov}(E[f(X_{1},X_{2})|\mathcal{F}],E[g(Y_{1},Y_{2})|\mathcal{F}]),$
where the conditional covariance $\mathrm{Cov}(U,V|\mathcal{F})$ of two random variables $U,V$ given the sigma field $\mathcal{F}$ is defined by
$\mathrm{Cov}(U,V|\mathcal{F})=E[UV|\mathcal{F}]-E[U|\mathcal{F}]E[V|\mathcal{F}].$
To simplify the notation, let
$I_{1}=E[\mathrm{Cov}(f(X_{1},X_{2}),g(Y_{1},Y_{2})|\mathcal{F})],I_{2}=\mathrm{Cov}(E[f(X_{1},X_{2})|\mathcal{F}],E[g(Y_{1},Y_{2})|\mathcal{F}]).$
Because $X_{2}$ and $Y_{1}$ are conditionally independent given $X_{1}$, $E[f(X_{1},X_{2})|\mathcal{F}]$ is a function of $X_{1}$. Similarly, $E[g(Y_{1},Y_{2})|\mathcal{F}]$ is a function of $Y_{1}$. Then, by definition of maximal correlation coefficients, we have that
$|I_{2}|\leq R(X_{1},Y_{1})\sqrt{\operatorname{Var}(E[f(X_{1},X_{2}))|\mathcal{F}])\operatorname{Var}(E[g(Y_{1},Y_{2})|\mathcal{F}])}.$
Next, we look for the upper bound of $I_{1}$. By definition of maximal correlation coefficients and independence between $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$, we have that
$|\operatorname{Cov}(f(X_{1},X_{2}),g(Y_{1},Y_{2})|\mathcal{F})|\leq R(X_{2},Y_{2})\sqrt{\operatorname{Var}(f(X_{1},X_{2})|\mathcal{F})\operatorname{Var}(g(Y_{1},Y_{2})|\mathcal{F})}.$
By taking the expectation on both sides and using Hölder’s inequality, we obtain that
$|I_{1}|$ $\leq E\left[R(X_{2},Y_{2})\sqrt{\operatorname{Var}(f(X_{1},X_{2})|\mathcal{F})\operatorname{Var}(g(Y_{1},Y_{2})|\mathcal{F})}\right]$
$\leq R(X_{2},Y_{2})\sqrt{E[\operatorname{Var}(f(X_{1},X_{2})|\mathcal{F})]E[\operatorname{Var}(g(Y_{1},Y_{2})|\mathcal{F})]}.$
For simplicity of notation, let
$A_{f}^{X}=E[\operatorname{Var}(f(X_{1},X_{2})|\mathcal{F})],A_{g}^{Y}=E[\operatorname{Var}(g(Y_{1},Y_{2})|\mathcal{F})],$
$B_{f}^{X}=\operatorname{Var}(E[f(X_{1},X_{2}))|\mathcal{F}]),B_{g}^{Y}=\operatorname{Var}(E[g(Y_{1},Y_{2})|\mathcal{F}]).$
Note that $A_{f}^{X}+B_{f}^{X}=\operatorname{Var}(f(X_{1},X_{2}))$, $A_{g}^{Y}+B_{g}^{Y}=\operatorname{Var}(g(Y_{1},Y_{2}))$. Then we have that
$|\operatorname{Cov}$ $(f(X_{1},X_{2}),g(Y_{1},Y_{2})|$
$\leq|I_{1}|+|I_{2}|$
$\leq R(X_{1},Y_{1})\sqrt{B_{f}^{X}B_{g}^{Y}}+R(X_{2},Y_{2})\sqrt{A_{f}^{X}A_{g}^{Y}}$
$\leq\max(R(X_{1},Y_{1}),R(X_{2},Y_{2}))(\sqrt{A_{f}^{X}A_{g}^{Y}}+\sqrt{B_{f}^{X}B_{g}^{Y}})$
$\leq\max(R(X_{1},Y_{1}),R(X_{2},Y_{2}))\sqrt{\sqrt{A_{f}^{X}}^{2}+\sqrt{B_{f}^{X}}^{2}}\sqrt{\sqrt{A_{g}^{Y}}^{2}+\sqrt{B_{g}^{Y}}^{2}}$
$=\max(R(X_{1},Y_{1}),R(X_{2},Y_{2}))\sqrt{\mathrm{Var}(f(X_{1},X_{2}))\mathrm{Var}(g(Y_{1},Y_{2}))}.$
Hence, $R((X_{1},X_{2}),(Y_{1},Y_{2}))\leq\max(R(X_{1},Y_{1}),R(X_{2},Y_{2}))$.
On the other hand, clearly, we have that
$R((X_{1},X_{2}),(Y_{1},Y_{2}))\geq\max(R(X_{1},Y_{1}),R(X_{2},Y_{2})).$
For the Csáki-Fisher identity, the independence between the random vectors $(X_{1},Y_{1})$ and $(X_{2},Y_{2})$ is necessary. We will show this by the following example:
Example 1.
Let $X_{1},Y_{1}$ and $\sigma$ be independent random variables with the same distribution such that $P(\sigma=-1)=P(\sigma=1)=1/2$. Since $X_{1}$ is independent of $Y_{1}$, we have $R(X_{1},Y_{1})=0$. Define
$X_{2}=\sigma Y_{1},Y_{2}=\sigma X_{1}.$
Then, we have the joint distribution of $X_{2},Y_{2}$, see Table 1:
$X_{2}\backslash Y_{2}$1 1
-1 1/4 1/4
1 1/4 1/4
Table 1: Joint distribution of $(X_{2},Y_{2})$
Immediately, we see that $X_{2}$ is independent of $Y_{2}$. Hence, we have that $R(X_{2},Y_{2})=0$. Note that $X_{1}X_{2}=Y_{1}Y_{2}=\sigma X_{1}Y_{1}$ and they are non-degenerate. Consequently, we have that
$R((X_{1},X_{2}),(Y_{1},Y_{2}))\geq R(X_{1}X_{2},Y_{1}Y_{2})=1>0=\max(R(X_{1},Y_{1}),R(X_{2},Y_{2})).$
2.4 Lower semi-continuity
The maximal correlation coefficient $R(X,Y)$ is fully determined by the joint distribution $\mu$ of $(X,Y)$. Therefore, for a joint distribution $\mu$ of $(X,Y)$, its maximal correlation coefficient $R(\mu)$ is well-defined by $R(\mu)=R(X,Y)$. Suppose that $(X,Y)$ take values in a complete separable metric space $S$ (i.e. Polish space) with the Borel-$\sigma$-field $\mathcal{S}$. Let $\mathcal{P}$ be the space of probability measures on $(S,\mathcal{S})$. We equip $\mathcal{P}$ with the Lévy-Prokhorov metric $\pi$, see [Bil99, Eq. (6.10)] for the definition of this metric. By [Bil99, Theorem 6.8], $(\mathcal{P},\pi)$ is Polish and weak convergence is equivalent to the convergence with respect to the metric $\pi$. In this section, we will show the lower semi-continuity of $\mu\mapsto R(\mu)$ with respect to the metric $\pi$. It is a generalization of [BDK05, Theorem 1].
Lemma 2.3.
Suppose that $(X,Y)$ takes values in a Polish space $(S,\mathcal{S})$ and let $\mu$ be the joint distribution of $(X,Y)$. As a function on the Polish space $(\mathcal{P},\pi)$, $\mu\mapsto R(\mu)=R(X,Y)$ is lower semi-continuous.
To prove Lemma 2.3, we need to use the classical result that $L^{p}$ functions can be approximated by bounded continuous functions. We give the precise statement in the following:
Lemma 2.4.
Consider a metric space S with Borel $\sigma$-field $\mathcal{S}$, a bounded measure $\mu$ on $(S,\mathcal{S})$, and a constant $p>0$. Then the bounded continuous functions on S are dense in $L^{p}(S,\mathcal{S},\mu)$. Thus, for any $f\in L^{p},$ there exist some bounded, continuous functions $f_{1},f_{2},\ldots:S\to\mathbb{R}$ with $\|f_{n}-f\|_{p}\to 0$.
The above lemma is precisely Lemma 1.37 in [Kal21].
Proof of Lemma 2.3.
Let $(X_{n},Y_{n})$ be a sequence of random variables with values in the measurable space $(S,\mathcal{S})$. Suppose that $(X_{n},Y_{n})$ converges weakly to $(X,Y)$ as $n\to+\infty$. It suffices to show that
$R(X,Y)\leq\liminf_{n\to+\infty}R(X_{n},Y_{n}).$
For any $\varepsilon>0$, by definition of the maximal correlation coefficient, there exists $\varphi$ and $\psi$ such that
$R(X,Y)\leq\rho(\varphi(X),\psi(Y))+\varepsilon$
and that $\operatorname{Var}(\varphi(X))>0$ and $\operatorname{Var}(\psi(Y))>0$. By Lemma 2.4 for $p=2$, for any $\delta>0$, there exist bounded continuous functions $f$ and $g$ such that
$E[(f(X)-\varphi(X))^{2}]<\delta,E[(g(Y)-\psi(Y))^{2}]<\delta.$
Since $\operatorname{Var}(\varphi(X))>0$ and $\operatorname{Var}(\psi(Y))>0$, for sufficiently small $\delta>0$, we have that $\operatorname{Var}(f(X))>0$, $\operatorname{Var}(g(Y))>0$ and $\rho(\varphi(X),\psi(Y))\leq\rho(f(X),g(Y))+\varepsilon$. By weak convergence of $(X_{n},Y_{n})$ towards $(X,Y)$, we have that
$\lim_{n\to\infty}\rho(f(X_{n}),g(Y_{n}))=\rho(f(X),g(Y)).$
By definition of maximal correlation coefficients,
$\rho(f(X_{n}),g(Y_{n}))\leq R(X_{n},Y_{n}).$
Hence, for any $\varepsilon>0$, we have that
$\liminf_{n\to\infty}R(X_{n},Y_{n})\geq\liminf_{n\to\infty}\rho(f(X_{n}),g(Y_{n}))=\rho(f(X),g(Y))\geq R(X,Y)-2\varepsilon.$
By taking $\varepsilon\to 0$, we obtain the desired result. ∎
Remark 2.1.
Later after we independently prove Lemma 2.3, we find a very similar statement in [KA16, Section II-A]. The idea of proofs are similar. However, there are several restrictions in their statement and proof. Firstly, they require that $X$ and $Y$ take values in finite state spaces. Secondly, the convergence of $(X_{n},Y_{n})$ towards $(X,Y)$ is assumed to be the convergence with respect to total variation distance instead of weak convergence.
3 Random walks
3.1 Proof of Theorem 1.1
In this subsection, we prove Theorem 1.1. As we have explained, the first equality is a direct consequence of Csáki-Fischer identity, see Theorem 2.2. To prove the second equality, for natural numbers $m\geq 1$, we define continuous time processes
$S^{(m)}_{t}=S_{\min(t,m)}\text{ and }T^{(m)}_{t}=T_{\min(t,m)},t\geq 0.$
Then, we have that
$R(S^{(m)},T^{(m)})=R((S_{n})_{n\leq m},(T_{n})_{n\leq m})=R(\xi_{1},\eta_{1}).$
As $m\to\infty$, the process $(S^{(m)},T^{(m)})$ converges to $(S,T)$ in the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$ by [Bil99, Theorem 16.1], where $D_{\mathbb{R}^{2}}[0,\infty)$ is the space of càdlàg functions $\omega:[0,\infty)\to\mathbb{R}^{2}$ endowed with the Skorokhod topology. Hence, by Lemma 2.3, we have that
$R(S,T)\leq\liminf_{m\to\infty}R(S^{(m)},T^{(m)})=R(\xi_{1},\eta_{1}).$
Finally, since $\xi_{1}=S_{1}-S_{0}$ is a measurable function of $S$ and $\eta_{1}=T_{1}-T_{0}$ is a measurable function of $T$, we have the opposite inequality
$R(\xi_{1},\eta_{1})\leq R(S,T).$
3.2 New proof of (3)
In this subsection, we give a new proof of (3) as a consequence of Theorem 1.1, the central limit theorem and Lemma 2.3.
Fix a Gaussian vector $(X,Y)$ such that the covariance is $\Sigma$ and the Pearson correlation coefficient is $r$.
Consider a random vector $\left(\xi,\eta\right)$ with the joint distribution in Table 2:
$\xi\backslash\eta$ $c$ $d$
$a$ $p_{ac}$ $p_{ad}$
$b$ $p_{bc}$ $p_{bd}$
Table 2: Joint distribution of $(\xi,\eta)$
Then, for a non-constant function $f$ on $\{a,b\}$ and a non-constant function $g$ on $\{c,d\}$, the Pearson correlation coefficient between $f(\xi)$ and $g(\eta)$ is
$\operatorname{sgn}(f(a)-f(b))\operatorname{sgn}(g(c)-g(d))\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}/\sqrt{p_{a}p_{b}p_{c}p_{d}}.$
Then, we have that $R(\xi,\eta)=\left|\rho(\xi,\eta)\right|=\left.\left|\det\begin{pmatrix}p_{ac}&p_{ad}\\ p_{bc}&p_{bd}\end{pmatrix}\right.\right|/\sqrt{p_{a}p_{b}p_{c}p_{d}}$. By proper choices of $a,b,c,d$ and the probability mass function $p$, we may assume that the covariance matrix of $(\xi,\eta)$ is precisely $\Sigma$, and the Pearson correlation coefficient equals to $r$.
Let $(\xi_{n},\eta_{n})_{n\geq 1}$ be i.i.d. random vectors such that $(\xi_{n},\eta_{n})$ has the same distribution as $(\xi,\eta)$. Define a two-dimensional random walk $(S,T)$ via (6). Then, by Theorem 1.1, we have that
$R\left(\frac{S_{m}-ES_{m}}{\sqrt{m}},\frac{T_{m}-ET_{m}}{\sqrt{m}}\right)\leq R(S,T)=R(\xi_{1},\eta_{1})=|r|.$
By the central limit theorem, as $m\to\infty$, $\left(\frac{S_{m}-ES_{m}}{\sqrt{m}},\frac{T_{m}-ET_{m}}{\sqrt{m}}\right)$ converges to $(X,Y)$ in distribution. By Lemma 2.3, we see that
$R(X,Y)\leq\liminf_{m\to\infty}R\left(\frac{S_{m}-ES_{m}}{\sqrt{m}},\frac{T_{m}-ET_{m}}{\sqrt{m}}\right)\leq|r|=|\rho(X,Y)|.$
On the other hand, by definition, we have that $R(X,Y)\geq|\rho(X,Y)|$. Thus, (3) is proved.
4 Lévy processes and stable distributions
4.1 Proof of Theorem 1.2
In this subsection, we prove Theorem 1.2. The proof crucially relies on the Lévy-Itô decomposition and the martingale representation theorem for Lévy processes. We will briefly present these classical results. We focus on two-dimensional Lévy processes. Without loss of generality, we may assume that the Lévy process is canonical. Take the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$ as the sample space $\Omega$, where $D_{\mathbb{R}^{2}}[0,\infty)$ is the space of càdlàg functions $\omega:[0,\infty)\to\mathbb{R}^{2}$ endowed with the Skorokhod topology and the corresponding Borel $\sigma$-field. For $\omega=(\omega_{t})_{t\geq 0}$, define $(X_{t},Y_{t})=\omega_{t}$ for $t\geq 0$. Let $P$ be a probability measure on $\Omega$ such that $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional Lévy process. Let
$\Delta X_{t}=X_{t}-\lim_{s\uparrow t}X_{s}$
be the jump of $(X_{t})_{t\geq 0}$ at time $t$. Similarly, we define $\Delta Y_{t}$. Then, the set $\{t\geq 0:\Delta X_{t}\neq 0\text{ or }\Delta Y_{t}\neq 0\}$ of jumping times is countable.
Theorem 4.1 (Lévy-Itô decomposition).
If $(X_{t},Y_{t})_{t\geq 0}$ is a two-dimensional Lévy process, then there exists $b=(b^{X},b^{Y})\in\mathbb{R}^{2}$, a two-dimensional Brownian motion $(B^{X}_{t},B^{Y}_{t})_{t\geq 0}$ with the covariance $\Sigma$ and a Poisson random measure $N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ with the intensity measure $\mathrm{d}t\nu(\mathrm{d}x,\mathrm{d}y)$ such that for $t\geq 0$,
$X_{t}=b^{X}t+B^{X}_{t}+\int_{0}^{t}\int_{|x|<1}x\widetilde{N}^{X}(\mathrm{d}s,\mathrm{d}x)+\int_{0}^{t}\int_{|x|\geq 1}xN^{X}(\mathrm{d}s,\mathrm{d}x),$ (14)
$Y_{t}=b^{Y}t+B^{Y}_{t}+\int_{0}^{t}\int_{|y|<1}y\widetilde{N}^{Y}(\mathrm{d}s,\mathrm{d}y)+\int_{0}^{t}\int_{|y|\geq 1}yN^{Y}(\mathrm{d}s,\mathrm{d}y),$
where
$N^{X}(\mathrm{d}t,\mathrm{d}x)=1_{x\neq 0}\int_{\mathbb{R}}N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y),N^{Y}(\mathrm{d}t,\mathrm{d}y)=1_{y\neq 0}\int_{\mathbb{R}}N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y),$
$\nu^{X}(\mathrm{d}x)=1_{x\neq 0}\int_{\mathbb{R}}\nu(\mathrm{d}x,\mathrm{d}y),\nu^{Y}(\mathrm{d}y)=1_{y\neq 0}\int_{\mathbb{R}}\nu(\mathrm{d}x,\mathrm{d}y)$
and
$\widetilde{N}^{X}(\mathrm{d}t,\mathrm{d}x)=N^{X}(\mathrm{d}t,\mathrm{d}x)-\mathrm{d}t\nu^{X}(\mathrm{d}x),\widetilde{N}^{Y}(\mathrm{d}t,\mathrm{d}y)=N^{Y}(\mathrm{d}t,\mathrm{d}y)-\mathrm{d}t\nu^{Y}(\mathrm{d}y).$
Moreover, $(B_{t})_{t\geq 0}$ is independent of $N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ and
$N(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)=\sum_{(s,\Delta X_{s},\Delta Y_{s}):(\Delta X_{s},\Delta Y_{s})\neq(0,0)}\delta_{(s,\Delta X_{s},\Delta Y_{s})}(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y),$
where $\delta_{z_{0}}(\mathrm{d}z)$ is the Dirac measure at $z_{0}\in[0,\infty)\times(\mathbb{R}^{2}\setminus\{0\})$.
We refer to [App09, Theorem 2.4.16] for a proof. Note that the random measure $N^{X}(\mathrm{d}t,\mathrm{d}x)$ is a measurable function of $(X_{t})_{t\geq 0}$. Here, the space of $\sigma$-finite measures is endowed with the weak topology. Moreover, almost surely, $t\mapsto\int_{0}^{t}\int_{|x|\geq 1}xN^{X}(\mathrm{d}s,\mathrm{d}x)$ is a random element in $D[0,\infty)$. Hence, the process $t\mapsto\int_{0}^{t}\int_{|x|\geq 1}xN^{X}(\mathrm{d}s,\mathrm{d}x)$ in $D[0,\infty)$ can be realized as a measurable function of $(X_{t})_{t\geq 0}$ after a modification up to a $P$-null set. For $n\geq 1$, define
$Z^{(n)}_{t}=\int_{0}^{t}\int_{1/n<|x|<1}x\widetilde{N}^{X}(\mathrm{d}s,\mathrm{d}x).$
Then, each $(Z^{(n)}_{t})_{t\geq 0}$ is a càdlàg martingale, and can be viewed as measurable function of $(X_{t})_{t\geq 0}$ with values in $D[0,\infty)$. Let
$Z_{t}=\int_{0}^{t}\int_{|x|<1}x\widetilde{N}^{X}(\mathrm{d}s,\mathrm{d}x).$
The process $(Z_{t})_{t\geq 0}$ is viewed as a càdlàg martingale, which is the limit of $(Z^{(n)}_{t})_{t\geq 0}$. To be more precise, it is proved that for fixed $t\geq 0$, $Z^{(n)}_{t}$ converges to $Z_{t}$ in $L^{2}$. Hence, by using Doob’s maximal inequality and taking subsequences if necessary, there exists a subsequence $Z^{(n_{k})}$ such that $Z^{(n_{k})}$ converges almost surely to $Z$ in the Skorokhod space $D[0,\infty)$. Hence, after a modification up to a $P$-null set, $(Z_{t})_{t\geq 0}$ is a measurable function of $(X_{t})_{t\geq 0}$. Then, by (14), the Brownian motion $(B^{X}_{t})_{t\geq 0}$ is also a measurable function of $(X_{t})_{t\geq 0}$. Similarly, $(B^{Y}_{t})_{t\geq 0}$ and $N^{Y}(\mathrm{d}t,\mathrm{d}y)$ are measurable functions of $(Y_{t})_{t\geq 0}$.
Next, we introduce the martingale representation. Fix $T>0$. Let $\mathcal{F}^{X}_{T}$ be the augmentation of $\sigma(X_{t},0\leq t\leq T)$.
Theorem 4.2 (The Itô representation).
Let $F$ be a square integrable $\mathcal{F}^{X}_{T}$-measurable random variable. Then, there exist square-integrable predictable processes $\varphi^{X}$ and $\psi^{X}$ such that almost surely,
$F$ $=E(F)+\int_{0}^{T}\varphi^{X}(t)\,\mathrm{d}B^{X}_{t}+\int_{0}^{T}\int_{\mathbb{R}\setminus\{0\}}\psi^{X}(t,x)\widetilde{N}^{X}(\mathrm{d}t,\mathrm{d}x)$
$=E(F)+\int_{0}^{T}\varphi^{X}(t)\,\mathrm{d}B^{X}_{t}+\int_{0}^{T}\int_{\mathbb{R}\setminus\{0\}}\int_{\mathbb{R}}\psi^{X}(t,x)\widetilde{N}(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y).$
Clearly, similar results hold for $G\in L^{2}(\Omega,\mathcal{F}^{Y}_{T},P)$. We refer to [App09, Section 5.3] for details. One way to prove the representation theorem is to use Wiener-Lévy chaos decomposition, see [Itô56, Theorem 2] and [Kun04, Theorem 1.1].
Now, let’s start to prove Theorem 1.2. Since $(B^{X}_{t})_{t\geq 0}$ is a measurable function of $(X_{t})_{t\geq 0}$, $B^{X}_{1}$ is a square-integrable measurable function of $(X_{t})_{t\geq 0}$. Similarly, $B^{Y}_{1}$ is a square-integrable measurable function of $(Y_{t})_{t\geq 0}$. Hence, we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\max(\rho(B^{X}_{1},B^{Y}_{1}),\rho(-B^{X}_{1},B^{Y}_{1}))=|\rho|.$ (15)
Next, by definition of $\mathrm{Op}(\nu)$, for each small $\varepsilon>0$, there exists measurable functions $\varphi(x)$ and $\psi(y)$ such that $\varphi(0)=\psi(0)=0$, $\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)>0$ and that
$\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)\geq(\mathrm{Op}(\nu)-\varepsilon)\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}.$
Let
$F$ $=\int_{0}^{1}\int_{\mathbb{R}\setminus\{0\}}\varphi(x)\widetilde{N}^{X}(\mathrm{d}t,\mathrm{d}x)$
$=\lim_{n\to\infty}\sum_{(t,\Delta X_{t}):t\leq 1,|\Delta X_{t}|>1/n}\varphi(\Delta X_{t})-\int_{\mathbb{R}^{2}}1_{|x|>1/n}\varphi(x)\,\nu(\mathrm{d}x,\mathrm{d}y),$
$G$ $=\int_{0}^{1}\int_{\mathbb{R}\setminus\{0\}}\psi(y)\widetilde{N}^{Y}(\mathrm{d}t,\mathrm{d}y)$
$=\lim_{n\to\infty}\sum_{(t,\Delta Y_{t}):t\leq 1,|\Delta Y_{t}|>1/n}\psi(\Delta Y_{t})-\int_{\mathbb{R}^{2}}1_{|y|>1/n}\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y).$
Then, $F$ is a square-integrable measurable function of $(X_{t})_{t\geq 0}$ and $G$ is a square integrable measurable function of $(Y_{t})_{t\geq 0}$. Moreover, we have that $E(F)=E(G)=0$ and that
$\operatorname{Var}(F)$ $=\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y),$
$\operatorname{Var}(G)$ $=\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y),$
$\operatorname{Cov}(F,G)$ $=\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y).$
Hence, we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\rho(F,G)\geq\mathrm{Op}(\nu)-\varepsilon.$
Since $\varepsilon>0$ is arbitrary, we see that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\mathrm{Op}(\nu).$ (16)
Hence, by (15) and (16), we have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\max(|\rho|,\mathrm{Op}(\nu))$.
To prove the reverse inequality $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\max(|\rho|,\mathrm{Op}(\nu))$, we fix $T>0$ and non-degenerate square-integrable $F=F((X_{t})_{0\leq t\leq T})$ and $G=G((Y_{t})_{0\leq t\leq T})$, where $F(\cdot)$ and $G(\cdot)$ are measurable functions. By [Bil99, Theorem 12.5], the Borel-$\sigma$-field $\mathcal{D}$ of the Skorokhod space $D[0,T]$ is equal to the sigma field generated by finite-dimensional cylinders. Hence, we have that $F\in\sigma(X_{t},0\leq t\leq T)\subset\mathcal{F}^{X}_{T}$ and $G\in\sigma(Y_{t},0\leq t\leq T)\subset\mathcal{F}^{Y}_{T}$. By Theorem 4.2, we find square-integrable predictable processes $\varphi^{X}$, $\psi^{X}$, $\varphi^{Y}$ and $\psi^{Y}$ such that $F=E(F)+F_{B}+F_{N}$ and $G=E(G)+G_{B}+G_{N}$ with probability $1$, where
$F_{B}=\int_{0}^{T}\varphi^{X}(t)\,\mathrm{d}B^{X}_{t},$ (17)
$G_{B}=\int_{0}^{T}\varphi^{Y}(t)\,\mathrm{d}B^{Y}_{t},$
$F_{N}=\int_{0}^{T}\int_{\mathbb{R}\setminus\{0\}}\psi^{X}(t,x)\widetilde{N}^{X}(\mathrm{d}t,\mathrm{d}x)=\int_{0}^{T}\int_{\mathbb{R}\setminus\{0\}}\int_{\mathbb{R}}\psi^{X}(t,x)\widetilde{N}(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y),$
$G_{N}=\int_{0}^{T}\int_{\mathbb{R}\setminus\{0\}}\psi^{Y}(t,y)\widetilde{N}^{Y}(\mathrm{d}t,\mathrm{d}y)=\int_{0}^{T}\int_{\mathbb{R}}\int_{\mathbb{R}\setminus\{0\}}\psi^{Y}(t,y)\widetilde{N}(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y).$
Claim 4.3.
We have that $E(F_{B}F_{N})=E(F_{B}G_{N})=E(G_{B}F_{N})=E(G_{B}G_{N})=0$.
Proof.
We only provide the proof for $E(F_{B}G_{N})=0$ as the other equations can be proved in a similar way. We define a martingale-valued measure $M$ as follows: for $t\geq 0$ and $A\subset\mathbb{R}^{2}$, we define
$M(t,A)=\int_{0}^{t}\int_{A\setminus\{0\}}\widetilde{N}(\mathrm{d}s,\mathrm{d}x,\mathrm{d}y)+B^{X}_{t}\delta_{0}(A).$
Let $\mu(t,A)=E(M(t,A)^{2})$. Then, by independence between $\widetilde{N}$ and $B^{X}$, we get that
$\mu(t,A)=t\nu(A\setminus\{0\})+t\Sigma_{11}\delta_{0}(A).$
Let $f(t,x,y)=1_{x=y=0}\varphi^{X}(t)$ and $g(t,x,y)=1_{y\neq 0}\psi^{Y}(t,y)$. Then, $F_{B}$ equals to the stochastic integration $\int_{0}^{T}f(t,x,y)M(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$ and $G_{N}$ equals to the stochastic integration $\int_{0}^{T}g(t,x,y)M(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)$. For the general theory of stochastic integration against certain type of martingale-valued measure, we refer to [App09, Chapter 4]. By Itô’s isometry for stochastic integrals (see e.g. [App09, Theorem 4.2.3 and Exercise 4.2.4]), we have that
$E(F_{B}G_{N})=\int_{0}^{T}\int_{\mathbb{R}^{2}}f(t,x,y)g(t,x,y)\,\mu(\mathrm{d}t,\mathrm{d}x,\mathrm{d}y)=0$
by definitions of $f$ and $g$. ∎
By Claim 4.3, we have that
[…]
By Itô’s isometry for stochastic integrals, we have that
$E(F_{B}^{2})$ $=\Sigma_{11}\int_{0}^{T}(\varphi^{X}(t))^{2}\,\mathrm{d}t,$
$E(G_{B}^{2})$ $=\Sigma_{22}\int_{0}^{T}(\varphi^{Y}(t))^{2}\,\mathrm{d}t,$
$E(F_{B}G_{B})$ $=\Sigma_{12}\int_{0}^{T}\varphi^{X}(t)\varphi^{Y}(t)\,\mathrm{d}t.$
By Cauchy-Schwarz inequality, we obtain that
$E(F_{B}G_{B})\leq|\rho|\sqrt{E(F_{B}^{2})E(G_{B}^{2})}.$ (19)
Next, we wish to prove that
$E(F_{N}G_{N})\leq\mathrm{Op}(\nu)\sqrt{E(F_{N}^{2})E(G_{N}^{2})}.$ (20)
Similarly, by Itô’s isometry for stochastic integrals against compensated Poisson random measure, we get that
$E(F_{N}^{2})$ $=\int_{0}^{T}\int_{\mathbb{R}^{2}}1_{x\neq 0}(\psi^{X}(t,x))^{2}\,\mathrm{d}t\nu(\mathrm{d}x,\mathrm{d}y),$
$E(G_{N}^{2})$ $=\int_{0}^{T}\int_{\mathbb{R}^{2}}1_{y\neq 0}(\psi^{Y}(t,y))^{2}\,\mathrm{d}t\nu(\mathrm{d}x,\mathrm{d}y),$
$E(F_{N}G_{N})$ $=\int_{0}^{T}\int_{\mathbb{R}^{2}}1_{x\neq 0,y\neq 0}\psi^{X}(t,x)\psi^{Y}(t,y)\,\mathrm{d}t\nu(\mathrm{d}x,\mathrm{d}y).$
By the definition of $\mathrm{Op}(\nu)$, we have that
$\int_{\mathbb{R}^{2}}1_{x\neq 0,y\neq 0}\psi^{X}(t,x)\psi^{Y}(t,y)\,\nu(\mathrm{d}x,\mathrm{d}y)\\ \leq\mathrm{Op}(\nu)\sqrt{\int_{\mathbb{R}^{2}}1_{x\neq 0}(\psi^{X}(t,x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}\sqrt{\int_{\mathbb{R}^{2}}1_{y\neq 0}(\psi^{Y}(t,y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}.$ (21)
Integrating both sides of (21) against $\mathrm{d}t$ and using Cauchy-Schwarz inequality, we prove (20). Finally, by (18), (19), (20) and Cauchy-Schwarz inequality, we have that
$\operatorname{Cov}(F,G)$ $\leq\max(|\rho|,\mathrm{Op}(\nu))\left(\sqrt{E(F_{B}^{2})E(G_{B}^{2})}+\sqrt{E(F_{N}^{2})E(G_{N}^{2})}\right)$
$\leq\max(|\rho|,\mathrm{Op}(\nu))\sqrt{E(F_{B}^{2})+E(F_{N}^{2})}\sqrt{E(G_{B}^{2})+E(G_{N}^{2})}$
$=\max(|\rho|,\mathrm{Op}(\nu))\sqrt{\operatorname{Var}(F)}\sqrt{\operatorname{Var}(G)}.$
Since the above inequality holds for all proper $F=F((X_{t})_{0\leq t\leq T})$ and $G=G((Y_{t})_{0\leq t\leq T})$, we get that
$R((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})\leq\max(|\rho|,\mathrm{Op}(\nu)).$
As $T\to\infty$, $((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})$ converges to $((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$, by Lemma 2.3 (lower-semi continuity), we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\liminf_{T\to\infty}R((X_{t})_{0\leq t\leq T},(Y_{t})_{0\leq t\leq T})\leq\max(|\rho|,\mathrm{Op}(\nu))$
and the proof is complete.
4.2 Proof of Theorem 1.3
We will prove Theorem 1.3 in this subsection. Firstly, we briefly explain the ideas in the following: We have found the maximal correlation coefficients for two dimensional Lévy processes in Theorem 1.2. Then, we apply Theorem 1.2 to two dimensional $\alpha$-stable processes $(X_{t},Y_{t})_{t\geq 0}$ with $20<\alpha<$ 2. The Brownian part vanishes. Hence, it suffices to calculate $\mathrm{Op}(\nu)$. For $\alpha$-stable processes, $\nu$ takes the special form (9). When $\tau(\mathrm{d}\theta)$ has a density $\tau(\theta)$ with respect to the Lebesgue measure $\mathrm{d}\theta$, the Lévy measure $\nu(\mathrm{d}x,\mathrm{d}y)$ also has a density $\nu(x,y)$. Moreover, if $\tau(-\theta)=\tau(\theta)$, we have $\nu(-x,-y)=\nu(x,y)$. In this case, we can define a homogeneous kernel $K(x,y)=\frac{\nu(x,y)}{\sqrt{\nu_{X}(x)\nu_{Y}(y)}}$ of degree $-1$, where $\nu_{X}(x)=\int_{\mathbb{R}}\nu(x,y)\,\mathrm{d}y$, $\nu_{Y}(y)=\int_{\mathbb{R}}\nu(x,y)\,\mathrm{d}x$ and $K(\lambda x,\lambda y)=|\lambda|^{-1}K(x,y)$ for all $x,y\in\mathbb{R}$ and $\lambda\neq 0$. The kernel $K(x,y)$ induces a linear operator $K$ on $L^{2}$ by $K\psi(x):=\int_{\mathbb{R}}K(x,y)\psi(y)\,\mathrm{d}y$. Then, the constant $\mathrm{Op}(\nu)$ is just the operator norm $||K||=\sup_{||\psi||_{2}>0}\frac{||K\psi||_{2}}{||\psi||_{2}}$. Here, “$\mathrm{Op}$” is short for the word “operator”. The determination of norm $||K||$ of homogenous kernel of degree $-1$ is a classical problem under the name Hilbert-Hardy inequality, see [HLP88, Theorem 319] and [Yan12, Theorem 42.8]. However, in our case, $\tau(\mathrm{d}\theta)$ is not necessarily absolutely continuous with respect to the Lebesgue measure. Moreover, even it is possible to define the kernel $K(x,y)$ in certain cases, due to the absence of the symmetry $\tau(-\theta)=\tau(\theta)$, $K(x,y)$ is not homogeneous. Instead, $K(x,y)$ is only positively homogeneous, i.e. $K(\lambda x,\lambda y)=\lambda^{-1}K(x,y)$ for $x,y\in\mathbb{R}$ and $\lambda>0$. The difference between homogeneity and positive homogeneity results in significant different expressions of $||K||$. We are not aware of existing results for $\mathrm{Op}(\nu)$. Hence, we calculate $\mathrm{Op}(\nu)$ in Lemma 4.4 by an adaptation of the argument leading to the Hilbert-Hardy inequalities. In this way, we successively find the maximal correlation coefficients $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ for $\alpha$-stable processes. As $\alpha$-stable distributions are marginal distributions of $\alpha$-stable processes, we have the upper bound $R(X_{1},Y_{1})\leq R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. The reverse inequality is a consequence of the convergence of stable random walks towards stable processes, Theorem 1.1 and Lemma 2.3. By the way, we could not generalize Theorem 1.3 to general infinitely divisible distributions. Indeed, it is possible to have $R(X_{1},Y_{1})<R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ for general two dimensional Lévy processes.
Lemma 4.4.
Suppose that the Lévy measure has the form (9). Let
$C_{++}=\int_{0}^{\frac{\pi}{2}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),\quad$ $C_{+-}=\int_{\frac{3\pi}{2}}^{2\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),$
$C_{-+}=\int_{\frac{\pi}{2}}^{\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),\quad$ $C_{--}=\int_{\pi}^{\frac{3\pi}{2}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta),$
and
$D^{x}_{+}=\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}|\cos\theta|^{\alpha}\,\tau(\mathrm{d}\theta),\quad$ $D^{x}_{-}=\int_{\frac{\pi}{2}}^{\frac{3\pi}{2}}|\cos\theta|^{\alpha}\,\tau(\mathrm{d}\theta),$
$D^{y}_{+}=\int_{0}^{\pi}|\sin\theta|^{\alpha}\,\tau(\mathrm{d}\theta),\quad$ $D^{y}_{-}=\int_{\pi}^{2\pi}|\sin\theta|^{\alpha}\,\tau(\mathrm{d}\theta).$
Then, we have that
$\mathrm{Op}(\nu)=\left\|\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}\right\|_{2}$
with the convention that $/0=0$ 0, where $\|\cdot\|_{2}$ means the spectral norm.
Proof.
Using the polar coordinates, by (9), we have that
$L:=\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)=\int_{0}^{\infty}\int_{\mathbb{R}/(2\pi\mathbb{Z})}\varphi(r\cos\theta)\psi(r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta).$
Since we require that $\varphi(0)=\psi(0)=0$, we may assume that $\cos\theta\neq 0$ and $\sin\theta\neq 0$ in the above integral. Write
$L=L_{++}+L_{-+}+L_{--}+L_{+-},$
where
$L_{++}=\int_{0}^{\infty}\int_{0}^{\frac{\pi}{2}}\varphi(r\cos\theta)\psi(r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta),$
$L_{-+}=\int_{0}^{\infty}\int_{\frac{\pi}{2}}^{\pi}\varphi(r\cos\theta)\psi(r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta),$
$L_{--}=\int_{0}^{\infty}\int_{\pi}^{\frac{3\pi}{2}}\varphi(r\cos\theta)\psi(r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta),$
$L_{+-}=\int_{0}^{\infty}\int_{\frac{3\pi}{2}}^{2\pi}\varphi(r\cos\theta)\psi(r\sin\theta)\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta).$
By writing the integrand as the product of
$\varphi(r\cos\theta)r^{-\frac{1+\alpha}{2}}|\tan\theta|^{\frac{\alpha}{4}}\text{ and }\psi(r\sin\theta)r^{-\frac{1+\alpha}{2}}|\cot\theta|^{\frac{\alpha}{4}},$
using Cauchy-Schwarz inequality, we obtain that
$L_{++}\leq\sqrt{I^{\varphi}_{++}I^{\psi}_{++}},$
where
$I^{\varphi}_{++}=\int_{0}^{\infty}\int_{0}^{\frac{\pi}{2}}(\varphi(r\cos\theta))^{2}\frac{1}{r^{1+\alpha}}|\tan\theta|^{\frac{\alpha}{2}}\,\mathrm{d}r\tau(\mathrm{d}\theta)$
and
$I^{\psi}_{++}=\int_{0}^{\infty}\int_{0}^{\frac{\pi}{2}}(\psi(r\sin\theta))^{2}\frac{1}{r^{1+\alpha}}|\cot\theta|^{\frac{\alpha}{2}}\,\mathrm{d}r\tau(\mathrm{d}\theta).$
By performing the change of variable $x=r\cos\theta$, we get that
$I^{\varphi}_{++}=\int_{0}^{\frac{\pi}{2}}\int_{0}^{\infty}(\varphi(x))^{2}\frac{1}{|x|^{1+\alpha}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}}\,\tau(\mathrm{d}\theta)\mathrm{d}x=C_{++}F_{+},$
where
$F_{+}=\int_{0}^{\infty}(\varphi(x))^{2}\frac{1}{|x|^{1+\alpha}}\,\mathrm{d}x.$
Similarly, by performing the change of variable $y=r\sin\theta$, we get that
$I^{\psi}_{++}=C_{++}G_{+},$
where
$G_{+}=\int_{0}^{\infty}(\psi(y))^{2}\frac{1}{|y|^{1+\alpha}}\,\mathrm{d}y.$
Hence, we have that
$L_{++}\leq C_{++}\sqrt{F_{+}}\sqrt{G_{+}}.$ (22)
Similarly, we have that
$L_{+-}\leq C_{+-}\sqrt{F_{+}}\sqrt{G_{-}},\quad L_{-+}\leq C_{-+}\sqrt{F_{-}}\sqrt{G_{+}},\quad L_{--}\leq C_{--}\sqrt{F_{-}}\sqrt{G_{-}},$
where
$F_{-}=\int_{-\infty}^{0}(\varphi(x))^{2}\frac{1}{|x|^{1+\alpha}}\,\mathrm{d}x,\quad G_{-}=\int_{-\infty}^{0}(\psi(y))^{2}\frac{1}{|y|^{1+\alpha}}\,\mathrm{d}y.$
Hence, we have that
$\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)\leq$ $C_{++}\sqrt{F_{+}}\sqrt{G_{+}}+C_{+-}\sqrt{F_{+}}\sqrt{G_{-}}$ (23)
$+C_{-+}\sqrt{F_{-}}\sqrt{G_{+}}+C_{--}\sqrt{F_{-}}\sqrt{G_{-}}.$
Similarly, we find that
$\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{x}_{+}F_{+}+D^{x}_{-}F_{-},$ (24)
where
$D^{x}_{+}=\int_{-\frac{\pi}{2}}^{\frac{\pi}{2}}|\cos\theta|^{\alpha}\,\tau(\mathrm{d}\theta),\quad D^{x}_{-}=\int_{\frac{\pi}{2}}^{\frac{3\pi}{2}}|\cos\theta|^{\alpha}\,\tau(\mathrm{d}\theta),$
and that
$\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{y}_{+}G_{+}+D^{y}_{-}G_{-},$ (25)
where
$D^{y}_{+}=\int_{0}^{\pi}|\sin\theta|^{\alpha}\,\tau(\mathrm{d}\theta),\quad D^{y}_{-}=\int_{\pi}^{2\pi}|\sin\theta|^{\alpha}\,\tau(\mathrm{d}\theta).$
Let
$A=\begin{pmatrix}C_{++}/\sqrt{D^{x}_{+}D^{y}_{+}}&C_{+-}/\sqrt{D^{x}_{+}D^{y}_{-}}\\ C_{-+}/\sqrt{D^{x}_{-}D^{y}_{+}}&C_{--}/\sqrt{D^{x}_{-}D^{y}_{-}}\end{pmatrix}$
with the convention that $/0=0$ 0. Let $||A||_{2}$ be the spectral norm of $A$. Then by (23), (24) and (25), we have that
$\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)\leq||A||_{2}\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}.$
Hence, $\mathrm{Op}(\nu)\leq||A||_{2}$. It remains to prove the converse inequality $\mathrm{Op}(\nu)\geq||A||_{2}$. Without loss of generality, we may assume that $||A||_{2}>0$. Equivalently, we assume that $D^{x}_{+}+D^{x}_{-}>0$ and $D^{y}_{+}+D^{y}_{-}>0$. For this purpose, for $\varepsilon>0$, we set
$\varphi(x)=1_{x>1}|x|^{\frac{\alpha}{2}-\varepsilon}b^{\varphi}_{+}+1_{x<-1}|x|^{\frac{\alpha}{2}-\varepsilon}b^{\varphi}_{-},\quad\psi(y)=1_{y>1}|y|^{\frac{\alpha}{2}-\varepsilon}b^{\psi}_{+}+1_{y<-1}|y|^{\frac{\alpha}{2}-\varepsilon}b^{\psi}_{-},$
where the positive constants $b^{\varphi}_{+}$, $b^{\varphi}_{-}$, $b^{\psi}_{+}$ and $b^{\psi}_{-}$ will be chosen later. Then, we have that
$L_{++}$ $=b^{\varphi}_{+}b^{\psi}_{+}\int_{0}^{\infty}\int_{0}^{\frac{\pi}{2}}1_{r\cos\theta>1,r\sin\theta>1}|r\cos\theta|^{\frac{\alpha}{2}-\varepsilon}|r\sin\theta|^{\frac{\alpha}{2}-\varepsilon}\frac{1}{r^{1+\alpha}}\,\mathrm{d}r\tau(\mathrm{d}\theta)$
$=b^{\varphi}_{+}b^{\psi}_{+}\int_{0}^{\infty}\int_{0}^{\frac{\pi}{2}}1_{r\cos\theta>1,r\sin\theta>1}|\cos\theta\sin\theta|^{\frac{\alpha}{2}-\varepsilon}\frac{1}{r^{1+2\varepsilon}}\,\mathrm{d}r\tau(\mathrm{d}\theta)$
$=\frac{1}{2\varepsilon}b^{\varphi}_{+}b^{\psi}_{+}\int_{0}^{\frac{\pi}{2}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}-\varepsilon}(\min(|\cos\theta|,|\sin\theta|))^{2\varepsilon}\,\tau(\mathrm{d}\theta).$
Similarly, we have that
$L_{+-}=\frac{1}{2\varepsilon}b^{\varphi}_{+}b^{\psi}_{-}\int_{\frac{3\pi}{2}}^{2\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}-\varepsilon}(\min(|\cos\theta|,|\sin\theta|))^{2\varepsilon}\,\tau(\mathrm{d}\theta),$
$L_{-+}=\frac{1}{2\varepsilon}b^{\varphi}_{-}b^{\psi}_{+}\int_{\frac{\pi}{2}}^{\pi}|\cos\theta\sin\theta|^{\frac{\alpha}{2}-\varepsilon}(\min(|\cos\theta|,|\sin\theta|))^{2\varepsilon}\,\tau(\mathrm{d}\theta),$
and
$L_{--}=\frac{1}{2\varepsilon}b^{\varphi}_{-}b^{\psi}_{-}\int_{\pi}^{\frac{3\pi}{2}}|\cos\theta\sin\theta|^{\frac{\alpha}{2}-\varepsilon}(\min(|\cos\theta|,|\sin\theta|))^{2\varepsilon}\,\tau(\mathrm{d}\theta).$
Moreover, we find that
$\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=D^{x}_{+}F_{+}+D^{x}_{-}F_{-}=\frac{1}{2\varepsilon}D^{x}_{+}(b^{\varphi}_{+})^{2}+\frac{1}{2\varepsilon}D^{x}_{-}(b^{\varphi}_{-})^{2}.$
and
$\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)=\frac{1}{2\varepsilon}D^{y}_{+}(b^{\psi}_{+})^{2}+\frac{1}{2\varepsilon}D^{y}_{-}(b^{\psi}_{-})^{2}.$
We chose $b^{\varphi}_{+},b^{\varphi}_{-},b^{\psi}_{+}$ and $b^{\psi}_{-}$ such that
$(\sqrt{D^{x}_{+}}b^{\varphi}_{+},\sqrt{D^{x}_{-}}b^{\varphi}_{-})A\begin{pmatrix}\sqrt{D^{y}_{+}}b^{\psi}_{+}\\ \sqrt{D^{y}_{-}}b^{\psi}_{-}\end{pmatrix}\\ =||A||_{2}\sqrt{(D^{x}_{+}(b^{\varphi}_{+})^{2}+D^{x}_{-}(b^{\varphi}_{-})^{2})(D^{y}_{+}(b^{\psi}_{+})^{2}+D^{y}_{-}(b^{\psi}_{-})^{2})}.$
Then, we get that
$\mathrm{Op}(\nu)\geq\lim_{\varepsilon\to 0}\frac{\int_{\mathbb{R}^{2}}\varphi(x)\psi(y)\,\nu(\mathrm{d}x,\mathrm{d}y)}{\sqrt{\int_{\mathbb{R}^{2}}(\varphi(x))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)\int_{\mathbb{R}^{2}}(\psi(y))^{2}\,\nu(\mathrm{d}x,\mathrm{d}y)}}=||A||_{2}.$
At present, we have found the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$. Clearly, we have that
$R(X_{1},Y_{1})\leq R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0}).$
On the other hand, consider the random walk $(S_{n},T_{n})_{n\geq 0}$ such that its increment $(S_{n+1}-S_{n},T_{n+1}-T_{n})$ has the same distribution as $(X_{1},Y_{1})$. Then, by Theorem 1.1, we have that
$R((S_{n})_{n\geq 0},(T_{n})_{n\geq 0})=R(X_{1},Y_{1}).$
Since $X_{1}$ is stable, there exists $(c_{n},d_{n})\in\mathbb{R}$ such that $(S_{n}-c_{n},T_{n}-d_{n})$ has the same distribution as $n^{1/\alpha}(X_{1},Y_{1})$. Define
$(X^{(n)}_{t},Y^{(n)}_{t})=\left(\frac{S_{[nt]}-c_{[nt]}}{n^{1/\alpha}},\frac{T_{[nt]}-d_{[nt]}}{n^{1/\alpha}}\right).$
Then, $R(X^{(n)},Y^{(n)})=R(S,T)=R(X_{1},Y_{1})$. Moreover, by [Kal21, Theorem 23.14], the process $(X^{(n)}_{t},Y^{(n)}_{t})_{t\geq 0}$ converges towards $(X_{t},Y_{t})_{t\geq 0}$ in the Skorokhod space $D_{\mathbb{R}^{2}}[0,\infty)$. Hence, by Lemma 2.3, we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\leq\liminf_{n\to\infty}R((X^{(n)}_{t})_{t\geq 0},(Y^{(n)}_{t})_{t\geq 0})=R(X_{1},Y_{1}).$
Finally, we find that $R(X_{1},Y_{1})=R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})=\mathrm{Op}(\nu)$.
From the above argument, we have the following observation.
Remark 4.1.
Suppose that $(X,Y)$ is a stable random vector. Let $(\widetilde{X},\widetilde{Y})$ be in the domain of contraction of $(X,Y)$. Then, we have that
$R(\widetilde{X},\widetilde{Y})\geq R(X,Y),$
where $R(X,Y)$ is given in Theorem 1.3.
4.3 Examples
Firstly, we consider an example studied in [BDK05]. They proved the following theorem:
Theorem 4.5 (Bryc-Dembo-Kagan).
Let $X$ and $Z$ be independent copies of $\alpha$-stable random variables with $0<\alpha\leq 2$. Then for all $\lambda\geq 0$, we have
$R(X,X+\lambda Z)=\frac{1}{\sqrt{1+|\lambda|^{\alpha}}}.$ (26)
If $X$ and $Z$ are symmetric, then the above inequality (26) holds for $\lambda<0$.
For $\alpha=2$, $X$ and $Z$ must be Gaussian and the result goes back to Lancaster [Lan57]. The main contribution is the case that $20<\alpha<$ 2, $\lambda\geq 0$ and $X$ is $\alpha$-stable. The result for $\lambda<0$ and symmetric $\alpha$-stable random variables can be easily deduced from (26) by taking $-Z$ instead of $Z$. The restriction to positive $\lambda$ in (26) has a reason: In general, the expression for $\lambda<0$ is different. Indeed, we have the following result:
Proposition 4.6.
Let $X$ and $Z$ be independent copies of $\alpha$-stable random variables with $20<\alpha<$ 2. Then, the Lévy measure $\nu_{X}(x)\mathrm{d}x$ of $X$ has the following form:
$\nu_{X}(x)\mathrm{d}x=\frac{c_{-}}{|x|^{1+\alpha}}1_{x<0}\,\mathrm{d}x+\frac{c_{+}}{|x|^{1+\alpha}}1_{x>0}\,\mathrm{d}x.$
Then, for all $\lambda<0$, we have
$R(X,X+\lambda Z)=1/\sqrt{1+\frac{\min(c_{-},c_{+})}{\max(c_{-},c_{+})}|\lambda|^{\alpha}}.$ (27)
The proof is based on Theorem 1.3 and the fact that $(X,X+\lambda Z)$ is stable.
Proof.
By using characteristic functions and independence between $X$ and $Z$, we find that $(X,Y)=(X,X+\lambda Z)$ is $\alpha$-stable random vector with Lévy measure $\nu(\mathrm{d}x,\mathrm{d}y)$, where
$\nu(\mathrm{d}x,\mathrm{d}y)=\frac{1}{\lambda}\nu_{X}\left(\frac{y}{\lambda}\right)\delta_{0}(\mathrm{d}x)\mathrm{d}y+\nu_{X}(x)\mathrm{d}x\delta_{x}(\mathrm{d}y).$ (28)
If we write $\nu(\mathrm{d}x,\mathrm{d}y)$ in the form (9), then we have
$\tau(\mathrm{d}\theta)=c_{-}|\lambda|^{\alpha}\delta_{\frac{\pi}{2}}(\mathrm{d}\theta)+c_{+}|\lambda|^{\alpha}\delta_{\frac{3\pi}{2}}(\mathrm{d}\theta)+c_{+}\sqrt{2}^{\alpha}\delta_{\frac{\pi}{4}}(\mathrm{d}\theta)+c_{-}\sqrt{2}^{\alpha}\delta_{\frac{5\pi}{4}}(\mathrm{d}\theta).$
Recall the notations in the statement of Theorem 1.3. Then, in our case, we have that
$C_{++}=c_{+},C_{--}=c_{-},C_{+-}=C_{-+}=0,$
$D^{x}_{+}=c_{+},D^{x}_{-}=c_{-},D^{y}_{+}=c_{+}+c_{-}|\lambda|^{\alpha},D^{y}_{-}=c_{-}+c_{+}|\lambda|^{\alpha}.$
Hence, by Theorem 1.3, we have that
$R(X,Y)$ $=\left\|\begin{pmatrix}\sqrt{\frac{c_{+}}{c_{+}+c_{-}|\lambda|^{\alpha}}}&0\\ 0&\sqrt{\frac{c_{-}}{c_{-}+c_{+}|\lambda|^{\alpha}}}\end{pmatrix}\right\|_{2}$
$=\max\left(\sqrt{\frac{c_{+}}{c_{+}+c_{-}|\lambda|^{\alpha}}},\sqrt{\frac{c_{-}}{c_{-}+c_{+}|\lambda|^{\alpha}}}\right)$
$=1/\sqrt{1+\frac{\min(c_{-},c_{+})}{\max(c_{-},c_{+})}|\lambda|^{\alpha}}.$
Secondly, we consider several compound Poisson processes.
Example 2.
Consider $b=0$, $\Sigma=0$ and a probability measure $\nu(\mathrm{d}x,\mathrm{d}y)$ in (1). Suppose that
$\nu(\{(x,y)\in\mathbb{R}^{2}\setminus\{0\}:x=0\text{ or }y=0\})=0.$
Let $(X_{t},Y_{t})$ be the corresponding Lévy processes. Then, we have that
$R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})=\mathrm{Op}(\nu)=1.$
The reason is that the number of jumps $N$ of $(X_{t})_{t\geq 0}$ up to time $1$ is equal to that of $(Y_{t})_{t\geq 0}$. Hence, $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq R(N,N)=1$. So, for generic compound Poisson processes, the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$ is equal to $1$.
Finally, we show that the maximal correlation coefficient $R(X_{1},Y_{1})$ could be strictly smaller than $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$.
Example 3.
Let $(B_{t})_{t\geq 0}$ be a (one dimensional) Brownian motion. Let $(N_{t})_{t\geq 0}$ be a Poisson process of rate $1$ which is independent of $(B_{t})_{t\geq 0}$. Let $X_{t}=B_{t}+N_{t}$. Then, $(X_{t},N_{t})_{t\geq 0}$ is a two dimensional Lévy process. By Theorem 1.2, we have that $R((X_{t})_{t\geq 0},(N_{t})_{t\geq 0})\geq\mathrm{Op}(\nu)=R((N_{t})_{t\geq 0},(N_{t})_{t\geq 0})=1$. Since the maximal correlation coefficients lie in the interval $[0,1]$, we have that $R((X_{t})_{t\geq 0},(N_{t})_{t\geq 0})=1$. However, $R(X_{1},N_{1})<1$. We will show this by contradiction. Assume that $R(X_{1},N_{1})=1$. Let $f(x,n)\,\mathrm{d}x=P(X_{1}\in(x,x+\mathrm{d}x),N_{1}=n)$, i.e.
$f(x,n)=\frac{e^{-1}}{n!}\frac{1}{\sqrt{2\pi}}e^{-(x-n)^{2}/2}.$
Let $f_{X_{1}}(x)$ be the marginal density of $X_{1}$ and $f_{N_{1}}(n)=P(N_{1}=n)=\frac{e^{-1}}{n!}$. Define
$k(x,n)=\frac{f(x,n)}{f_{X_{1}}(x)f_{N_{1}}(n)}$
and
$C^{2}$ $=\int_{-\infty}^{\infty}\sum_{n=0}^{\infty}(k(x,n)-1)^{2}f_{X_{1}}(x)f_{N_{1}}(n)\,\mathrm{d}x$
$=\int_{-\infty}^{\infty}\sum_{n=0}^{\infty}\frac{f(x,n)^{2}}{f_{X_{1}}(x)f_{N_{1}}(n)}\,\mathrm{d}x-1$
$=\int_{-\infty}^{\infty}\sum_{n=0}^{\infty}f_{N_{1}|X_{1}}(n|x)f_{X_{1}|N_{1}}(x|n)\,\mathrm{d}x-1,$
where $f_{N_{1}|X_{1}}(n|x)=P(N_{1}=n|X_{1}=x)$ and $f_{X_{1}|N_{1}}(x|n)$ is the conditional density of $X_{1}$ given $N_{1}=n$. Here, $C$ is called mean square contingency of $X_{1}$ and $N_{1}$. We will show that $C<\infty$ as follows: Note that
$f_{X_{1}|N_{1}}(x|n)=\frac{1}{\sqrt{2\pi}}e^{-(x-n)^{2}/2}.$ (29)
Note that
$f_{N_{1}|X_{1}}(n|x)=\frac{f(x,n)}{f_{X_{1}}(x)}=\frac{1}{f_{X_{1}}(x)}P(N_{1}=n)\frac{1}{\sqrt{2\pi}}e^{-(x-n)^{2}/2},$
$f_{X_{1}}(x)=\sum_{n=0}^{\infty}f(x,n)=\sum_{n=0}^{\infty}P(N_{1}=n)\frac{1}{\sqrt{2\pi}}e^{-(x-n)^{2}/2}.$
For $x\leq 0$, we have that
$f_{X_{1}}(x)\geq P(N_{1}=0)\frac{1}{\sqrt{2\pi}}e^{-x^{2}/2}=\frac{e^{-1}}{\sqrt{2\pi}}e^{-x^{2}/2}.$
Hence, for $x\leq 0$, we have that
$f_{N_{1}|X_{1}}(n|x)\leq\frac{1}{n!}e^{-n^{2}/2-n|x|}.$ (30)
Hence, by (29) and (30), we see that
$\int_{-\infty}^{0}\sum_{n=0}^{\infty}f_{N_{1}|X_{1}}(n|x)f_{X_{1}|N_{1}}(x|n)\,\mathrm{d}x<\infty.$
It remains to prove that
$\int_{0}^{\infty}\sum_{n=0}^{\infty}f_{N_{1}|X_{1}}(n|x)f_{X_{1}|N_{1}}(x|n)\,\mathrm{d}x<\infty.$ (31)
Hence, we assume that $x>0$ in the following. For fixed $x\geq 0$, let
$g(x)=\max_{n\geq 0}f(x,n)=\max_{n\geq 0}P(N_{1}=n)\frac{1}{\sqrt{2\pi}}e^{-(x-n)^{2}/2}.$
Then, there exists $c>0$ such that for $x\geq e$, we have that
$g(x)\geq f(x,\lceil x-\log x\rceil)\geq\frac{e^{-1}}{\sqrt{2\pi}}\frac{1}{\lceil x-\log x\rceil!}e^{-(\log x)^{2}/2}\geq\frac{c}{\lceil x\rceil^{10}}\frac{1}{\lceil x\rceil!}e^{(\log\lceil x\rceil)^{2}/2},$
where $\lceil x\rceil$ is the least integer that is greater than or equal to $x$. Then,
$f_{X_{1}}(x)=\sum_{n=0}^{\infty}f(x,n)\geq\max_{n\geq 0}f(x,n)=g(x)$
and for $x\geq e$,
$f_{N_{1}|X_{1}}(n|x)\leq\frac{1}{g(x)}\frac{e^{-1}}{n!}\frac{1}{\sqrt{2\pi}}e^{-(x-n)^{2}/2}\leq\frac{1}{c}(x+1)^{10}\frac{\lceil x\rceil!}{n!}e^{-(x-n)^{2}/2-(\log\lceil x\rceil)^{2}/2}.$
Note that $\lceil x\rceil!/n!\leq 1$ for $n\geq\lceil x\rceil$ and $\lceil x\rceil!/n!\leq\lceil x\rceil^{\lceil x\rceil-n}$ for $0\leq n\leq\lceil x\rceil$. Hence, there exists $c>0$ and $C<\infty$ such that for $x>0$ and $|n-\lceil x\rceil|\leq\log\lceil x\rceil/10$,
$f_{N_{1}|X_{1}}(n|x)\leq Ce^{-c(\log\lceil x\rceil)^{2}}.$ (32)
Besides, we have the trivial bound
$f_{N_{1}|X_{1}}(n|x)=P(N_{1}=n|X_{1}=x)\leq 1\text{ for }|n-\lceil x\rceil|>\log\lceil x\rceil/10.$ (33)
Combining (29), (32) and (33), we get (31). Hence, the mean square contingency $C$ is finite. By [R5́9, Theorem 2], there exists non-degenerate $\varphi(X_{1})$ and $\psi(N_{1})$ such that $\rho(\varphi(X_{1}),\psi(N_{1}))=R(X_{1},N_{1})=1$. Therefore, there exist $c_{1}\neq 0,c_{2}\neq 0$ and $d\in\mathbb{R}$ such that $P(c_{1}\varphi(X_{1})+c_{2}\psi(N_{1})=d)=1$. However, given the value of $N_{1}$, $X_{1}$ could take any value in $\mathbb{R}$. If $P(c_{1}\varphi(X_{1})+c_{2}\psi(N_{1})=d)=1$, then $\varphi$ is constant almost everywhere, which contradicts with the non-degeneracy of $\varphi(X_{1})$. Finally, by contradiction, we prove that $R(X_{1},N_{1})<1$.
For compound Poisson processes, it is still possible that $R(X_{1},Y_{1})<R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$, see the following example:
Example 4.
Let $(M_{t})_{t\geq 0}$ and $(N_{t})_{t\geq 0}$ be two independent Poisson processes of rate $1$. Let $X_{t}=M_{t}-N_{t}$ and $Y_{t}=M_{t}$. Then, $(X_{t},Y_{t})_{t\geq 0}$ is a two dimensional compound Poisson process. We have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})\geq\mathrm{Op}(\nu)=1$. Since the maximal correlation coefficients is at most one, we have that $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})=1$. However, we find that $R(X_{1},Y_{1})<1$ by numerical methods although we don’t have a theoretical proof at present. Indeed, let $X^{(n)}_{1}=\min(\max(X_{1},-n),n)$ and $Y^{(n)}_{1}=\min(\max(Y_{1},-n),n)$. Since $X^{(n)}_{1}$ is a measurable function of $X_{1}$ and $Y^{(n)}_{1}$ is a measurable function of $Y_{1}$, we have that $R(X^{(n)}_{1},Y^{(n)}_{1})\leq R(X_{1},Y_{1})$. On the other hand, since $(X^{(n)}_{1},Y^{(n)}_{1})$ converges to $(X_{1},Y_{1})$ in distribution, by Lemma 2.3, we have that $R(X,Y)\leq\liminf_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1})$. Therefore, we must have
$R(X,Y)=\lim_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1}).$
Since $(X^{(n)}_{1},Y^{(n)}_{1})$ takes values in a finite set, the maximal correlation $R(X^{(n)}_{1},Y^{(n)}_{1})$ is given by certain eigenvalue of a finite matrix, which could be found by numerical methods. By numerical calculation, we found that $R(X,Y)=\lim_{n\to\infty}R(X^{(n)}_{1},Y^{(n)}_{1})$ is close to $0.6925$.
5 Analog and generalization of DKS inequality
5.1 Proof of Proposition 1.4
We prove Proposition 1.4 in this subsection.
Recall the definition of $(X_{1},X_{2},\ldots,X_{n})$ and $(Y_{1},Y_{2},\ldots,Y_{n})$.
Firstly, consider the case $n=1$. Let $X$ be a non-degenerate random variable. Let $B$ is a Bernoulli random variable independent of $X$. Assume that $P(B=1)=1-P(B=0)=p$. Define
$Y=\left\{\begin{array}[]{ll}X,&\text{if }B=1,\\ \partial,&\text{if }B=0.\end{array}\right.$
Lemma 5.1.
The maximum correlation coefficient between $X$ and $Y$ is equal to $\sqrt{p}$.
Proof.
To calculate the maximum correlation coefficient between $X$ and $Y$, we take two functions $f$ and $g$ such that $E[f(X)]=0$, $\operatorname{Var}[f(X)]<\infty$, $E[g(Y)]=0$ and $\operatorname{Var}[g(Y)]<\infty$. Then, we have that
$\operatorname{Var}[g(Y)]=E[(g(Y))^{2}]=E[E[(g(Y))^{2}|B]]=pE[(g(X))^{2}]+(1-p)(g(\partial))^{2}.$ (34)
We calculate the covariance between $f(X)$ and $g(Y)$ by taking the expectation conditionally on $B$:
$\operatorname{Cov}(f(X),g(Y))$ $=E[f(X)g(Y)]$
$=E[E[f(X)g(Y)|B]]$
$=pE[f(X)g(X)]+(1-p)g(\partial)E[f(X)]$
$=pE[f(X)g(X)],$
where the first and the last inequalities are due to $E[f(X)]=0$. Combining previous results, we get that
$\rho(f(X),g(Y))$ $=\frac{\operatorname{Cov}(f(X),g(Y))}{\sqrt{\operatorname{Var}(f(X))}\sqrt{\operatorname{Var}(g(Y))}}$
$=\frac{pE[f(X)g(X)]}{\sqrt{E[(f(X))^{2}]}\sqrt{(1-p)(g(\partial))^{2}+pE[(g(X))^{2}]}}$ (35)
$\leq\frac{pE[f(X)g(X)]}{\sqrt{E[(f(X))^{2}]}\sqrt{pE[(g(X))^{2}]}}$
$\leq\sqrt{p},$
where we use Cauchy-Schwarz inequality $(E[f(X)g(X)])^{2}\leq E[(f(X))^{2}]E[(g(X))^{2}]$ in the last step. Thus, we have proved that $R(X,Y)\leq\sqrt{p}$.
Finally, there exists $g$ such that $g(\partial)=0$, $E[g(X)|X\neq\partial]=0$ and $0<E[(g(X))^{2}|X\neq\partial]<\infty$. Take $f=g$. Then, $E[f(X)]=0$ and $0<\operatorname{Var}[f(X)]=E[(g(X))^{2}]<\infty$. Moreover, $E[g(Y)]=pE[g(X)]=0$ and $0<\operatorname{Var}[g(Y)]<\infty$ by (34). Then, by (5.1), we have that
$R(X,Y)\geq\rho(f(X),g(Y))=\sqrt{p}$
and the proof is complete. ∎
By Lemma 5.1, we have that
$R(X_{i},Y_{i})=\sqrt{p_{i}}.$
Since $(X_{1},Y_{1}),(X_{2},Y_{2}),\ldots,(X_{n},Y_{n})$ are independent, by Csáki-Fischer identity, we have
$R((X_{1},X_{2},\ldots,X_{n}),(Y_{1},Y_{2},\ldots,Y_{n}))$ $=\max\{R(X_{i},Y_{i}):i=1,2,\ldots,n\}$
$=\sqrt{\max(p_{1},p_{2},\ldots,p_{n})}.$
Remark 5.1.
In the above arguments, we need the condition that $\partial$ is out of $F$ to ensure that the distribution of $X$ conditionally on $X\neq\partial$ is non-degenerate, which guarantee the existence of $g$ such that $g(\partial)=0$, $E[g(X)|X\neq\partial]=0$ and $0<E[(g(X))^{2}|X\neq\partial]<\infty$. It is possible to allow $P(X_{i}=\partial)>0$ for some $i$. However, in such cases, we need to assume that the distribution of $X_{i}$ conditionally on $X_{i}\neq\partial$ is non-degenerate. Otherwise, the formula is no longer correct.
5.2 Proof of Theorem 1.6
We prove Theorem 1.6 in this subsection. The key ingredient is the analysis of variance (ANOVA) decomposition developed in [MB07, Appendix I]. For the convenience of readers, we briefly explain the ANOVA decomposition without proofs. Let $X_{1},X_{2},\ldots,X_{n}$ be independent random variables. Write $X=(X_{1},X_{2},\ldots,X_{n})$. Suppose $\psi:F^{n}\to\mathbb{R}$ belongs to $L^{2}$, i.e. $\psi$ is a measurable function such that $E\psi^{2}(X_{1},X_{2},\ldots,X_{n})<\infty$. For each $j\in[n]$, define $E_{j}\psi$ by
$E_{j}\psi(x_{1},x_{2},\ldots,x_{n})=E[\psi(X_{1},X_{2},\ldots,X_{n})|X_{i}=x_{i},\forall i\neq j].$
In particular, $E_{j}\psi$ does not depend on $x_{j}$. For a subset $t\subset[n]$, define the linear subspace
$\mathcal{H}_{t}=\{\psi\in L^{2}:E_{j}\psi=\psi 1_{j\notin t},\forall j\in[n]\}.$
In particular, for $\psi\in\mathcal{H}_{t}$, $\psi$ does not depend on $x_{j}$ for $j\notin t$. Then, $(\mathcal{H}_{t})_{t\subset[n]}$ are orthogonal. Denote by $\overline{E}_{t}$ the orthogonal projection from $L^{2}$ onto $\mathcal{H}_{t}$. In fact, $\overline{E}_{t}=\prod_{j\in t}(I-E_{j})\prod_{k\notin t}E_{k}$, where $I$ is the identity map. In particular, $\overline{E}_{\emptyset}$ equals to the usual expectation $E$. Then, we have the orthogonal decomposition
$\psi=\sum_{t\subset[n]}\overline{E}_{t}\psi.$
If $\psi$ depends only on $(x_{j})_{j\in s}$ for some $s\subset[n]$ and $t$ is not a subset of $s$, then $\overline{E}_{t}\psi=0$ by definition of $\overline{E}_{t}$ and the fact that $E_{j}\psi=\psi$ for $j\in t\setminus s$. Hence, if $\psi$ depends only on $(x_{j})_{j\in s}$, then we have that
$\psi=\sum_{t\subset s}\overline{E}_{t}\psi.$
Sometimes, we encounter the random variable $\overline{E}_{t}\psi(X)$. Since $\overline{E}_{t}\psi$ depends only on $(x_{j})_{j\in t}$, we may write $\overline{E}_{t}\psi(X_{t})$ instead of $\overline{E}_{t}\psi(X)$, where $X_{t}$ means the sub-vector $(X_{j})_{j\in t}$.
Then, we are ready for the proof of Theorem 1.6. For a vector $x=(x_{1},x_{2},\ldots,x_{n})$ and a subset $s\subset[n]$, we denote by $x_{s}$ the sub-vector $(x_{j})_{j\in s}$. In this way, we have $X_{S}=(X_{j})_{j\in S}$ and $X_{T}=(X_{j})_{j\in T}$. Since $\partial$ is out of $(F,\mathcal{F})$, we have that
$R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))=R((S,X_{S}),(T,X_{T})).$
Consider two functions $\varphi$ and $\psi$ such that $E\varphi(S,X_{S})=E\psi(T,X_{T})=0$, $\operatorname{Var}\varphi(S,X_{S})<\infty$ and $\operatorname{Var}\varphi(T,X_{T})<\infty$. For $s\subset[n]$, define a function $\varphi_{s}$ via
$\varphi_{s}(x_{s})=\varphi(s,x_{s}).$
Note that $\varphi_{s}$ depends only on $(x_{j})_{j\in s}$. Similarly, we define $\psi_{t}$ for $t\subset[n]$. Then, by independence between $(S,T)$ and $X$, we have that
$\operatorname{Var}\varphi(S,X_{S})$ $=E\varphi^{2}(S,X_{S})$
$=\sum_{s\subset[n]}P(S=s)E\varphi_{s}^{2}(X_{s})$
$=\sum_{u\subset[n]}\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u}),$ (36)
where
$\varphi_{s,u}=\overline{E}_{u}\varphi_{s}$
and that the last equality of (5.2) comes from the orthogonality in the ANOVA decomposition $\varphi_{s}=\sum_{u:u\subset s}\overline{E}_{u}\varphi_{s}$. Similarly, we have that
$\operatorname{Var}\psi(T,X_{T})=\sum_{u\subset[n]}\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u}),$ (37)
where
$\psi_{t,u}=\overline{E}_{u}\psi_{t}.$
Moreover, we have that
$\operatorname{Cov}(\varphi(S,X_{S}),\psi(T,X_{T}))$ $=E[\varphi(S,X_{S})\psi(T,X_{T})]$
$=\sum_{u\subset[n]}\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)E[\varphi_{s,u}(X_{u})\psi_{t,u}(X_{u})]$
$\leq\sum_{u\subset[n]}\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E\varphi_{s,u}^{2}(X_{u})}\sqrt{E\psi_{t,u}^{2}(X_{u})}.$ (38)
For $u\neq\emptyset$, we have that
$\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E\varphi_{s,u}^{2}(X_{u})}\sqrt{E\psi_{t,u}^{2}(X_{u})}\\ \leq r_{u}\sqrt{\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})},$ (39)
where $r_{u}$ is the best constant $r$ for the inequality
$\sum_{s,t}P(S=s,T=t)\alpha_{s}\beta_{t}\leq r\sqrt{\sum_{s}P(S=s)\alpha_{s}^{2}}\sqrt{\sum_{t}P(T=t)\beta_{t}^{2}}$ (40)
for all real $\alpha_{s}$ and $\beta_{t}$ such that $\alpha_{s}=0$ if $s$ does not contain $u$ and $\beta_{t}=0$ if $t$ does not contain $u$.
For $u=\emptyset$, $\varphi_{s,u}=E\varphi_{s}(X_{s})$ and $\psi_{t,u}=E\psi_{t}(X_{t})$. Since $E\varphi(S,X_{S})=E\psi(T,X_{T})=0$, we have that
$\sum_{s\subset[n]}P(S=s)E\varphi_{s}(X_{s})=\sum_{t\subset[n]}P(T=t)E\varphi_{t}(X_{t})=0.$
Define $f(s)=E\varphi_{s}(X_{s})$ and $g(t)=E\varphi_{t}(X_{t})$. Then, we have that $Ef(S)=Eg(T)=0$, $\operatorname{Var}f(S)<\infty$ and $\operatorname{Var}g(T)<\infty$. By definition of maximal correlation coefficients, we have that
$\operatorname{Cov}(f(S),g(T))\leq R(S,T)\sqrt{\operatorname{Var}f(S)}\sqrt{\operatorname{Var}g(T)}.$
For $u=\emptyset$, we have that $\varphi_{s,u}(X_{u})=f(s)$, $\psi_{t,u}(X_{u})=g(t)$ and
$\sum_{(s,t):s\cap t\supset u}P(S=s,T=t)\sqrt{E\varphi_{s,u}^{2}(X_{u})}\sqrt{E\psi_{t,u}^{2}(X_{u})}\\ \leq R(S,T)\sqrt{\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})}.$ (41)
By (5.2), (37), (5.2), (39), (41) and Cauchy-Schwarz inequality, we have that
[…]
×$\quad\times\sum_{u\subset[n]}\sqrt{\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})}$
$\leq\max(R(S,T),\max(r_{u}:u\neq\emptyset))$
×$\quad\times\sqrt{\sum_{u\subset[n]}\sum_{s:s\supset u}P(S=s)E\varphi_{s,u}^{2}(X_{u})}\sqrt{\sum_{u\subset[n]}\sum_{t:t\supset u}P(T=t)E\psi_{t,u}^{2}(X_{u})}$
$=\max(R(S,T),\max(r_{u}:u\neq\emptyset))\sqrt{\operatorname{Var}\varphi(S,X_{S})}\sqrt{\operatorname{Var}\psi(T,X_{T})}.$ (42)
As $\varphi$ and $\psi$ is arbitrary, we conclude that
$R((S,X_{S}),(T,X_{T}))\leq\max(R(S,T),\max(r_{u}:u\neq\emptyset)).$
By definition of $r_{u}$, for $u\subset\tilde{u}$, we have that $r_{u}\geq r_{\tilde{u}}$. Hence, we have that
$\max(r_{u}:u\neq\emptyset)=\max(r_{\{j\}}:j\in[n]).$
It remains to prove the converse inequality
$R((S,X_{S}),(T,X_{T}))\geq\max(R(S,T),\max(r_{\{j\}}:j\in[n])).$
For $j\in[n]$, we take
$\varphi(S,X_{S})=1_{j\in S}\alpha_{S}h(X_{j}),\quad\psi(T,X_{T})=1_{j\in T}\beta_{T}h(X_{j})$
for some measurable function $h$ such that $Eh(X_{j})=0$, $\operatorname{Var}h(X_{j})<\infty$ and for $\alpha_{s}$ and $\beta_{t}$ achieving the best constant in the definition of $r_{\{j\}}$ (see (40)). It is strait-forward to check that for this particular choices of functions, we have that
$\operatorname{Cov}(\varphi(S,X_{S}),\psi(T,X_{T}))=r_{\{j\}}\sqrt{\operatorname{Var}\varphi(S,X_{S})\operatorname{Var}\psi(T,X_{T})}.$
Hence, $R((S,X_{S}),(T,X_{T}))\geq r_{\{j\}}$ for each $j\in[n]$. On the other hand, it is clear that
$R((S,X_{S}),(T,X_{T}))\geq R(S,T).$
Hence, the converse inequality is proved.
When $S$ and $T$ are independent, we have $R(S,T)=0$. By writing $\widetilde{\alpha}_{s}=\sqrt{P(S=s)}\alpha_{s}$ and $\widetilde{\beta}_{t}=\sqrt{P(T=t)}\beta_{t}$, we see that $r_{\{j\}}$ is exactly the spectral radius of the rank-one matrix $(M_{st})_{s,t:j\in s\cap t}$, where $M_{st}=\sqrt{P(S=s)P(T=t)}$. Hence, we have that
$r_{\{j\}}=\sqrt{P(j\in S)P(j\in T)}.$
5.3 Proof of Theorem 1.7
We prove Theorem 1.7 in this subsection. We will use the alternative definition (2) of the maximal correlation coefficient, the mathematical induction and the following lemma.
Lemma 5.2.
Let $T$ be a uniform subset of $[n]=\{1,2,\ldots,n\}$ of size $n-1$. Given $T$, the random set $S$ is a uniform subset of $T$ of size $k$. Then, we have that
$R(S,T)=\sqrt{\frac{k}{(n-1)(n-k)}}.$
Proof.
Let $\varphi$ be a measurable function such that $E\varphi(T)=0$ and $\operatorname{Var}\varphi(T)<\infty$. Then, we have that
$E(\varphi(T)|S=s)=\sum_{i=1}^{n}\varphi(\{i\}^{c})P(T=\{i\}^{c}|S=s)$
and that
$E[(E(\varphi(T)|S)^{2}]$ $=\sum_{s\subset[n]}P(S=s)\left(\sum_{i=1}^{n}\varphi(\{i\}^{c})P(T=\{i\}^{c}|S=s)\right)^{2}$
$=\sum_{s\subset[n]}\sum_{i=1}^{n}\sum_{j=1}^{n}\varphi(\{i\}^{c})\varphi(\{j\}^{c})P(S=s)$
×$\quad\quad\quad\quad\quad\quad\times P(T=\{i\}^{c}|S=s)P(T=\{j\}^{c}|S=s)$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\varphi(\{i\}^{c})\varphi(\{j\}^{c})A_{ij},$
where
$A_{ij}$ $=\sum_{S\subset[n]}P(S=s)P(T=\{i\}^{c}|S=s)P(T=\{j\}^{c}|S=s)$
$=\frac{1}{(n-k)^{2}}P(i\notin S,j\notin S)$
=$=\left\{\begin{array}[]{ll}\frac{1}{(n-k)n},&i=j,\\ \frac{(n-k-1)}{n(n-1)(n-k)},&i\neq j.\end{array}\right.$
The only element outside of $T$ is uniformly distributed in $[n]$. Hence, the condition $E\varphi(T)=0$ and $\operatorname{Var}\varphi(T)<\infty$ is equivalent to
$\sum_{i=1}^{n}\varphi(\{i\}^{c})=0$
and the value $\varphi(\{i\}^{c})$ is finite for each $i\in[n]$. Hence, by (2), we have that
$R^{2}(S,T)=\sup\{\sum_{i,j=1}^{n}A_{ij}\varphi(\{i\}^{c})\varphi(\{j\}^{c}):\sum_{i=1}^{n}\varphi(\{i\}^{c})=0,\sum_{i=1}^{n}(\varphi(\{i\}^{c}))^{2}=n\}.$
Note that $A=(A_{ij})_{i,j\in[n]}$ is a real symmetric matrix with positive elements. Hence, the eigenvalues of $A$ are real. By Perron-Frobenius theorem, $A$ has a simple eigenvalue $1/n$ and the moduli of other eigenvalues are strictly smaller than $1/n$. The eigenvector associated with $1/n$ is the column vector $\bm{1}=(1,1,\ldots,1)^{T}$. And the eigenvectors associated with other eigenvalues are perpendicular to $\bm{1}$. Hence, we get that
$R^{2}(S,T)=n\lambda_{2}(A),$
where $\lambda_{2}(A)$ is the second largest eigenvalue of $A$. Note that
$A=\frac{k}{n(n-1)(n-k)}I+\frac{n-k-1}{n(n-1)(n-k)}\bm{1}\bm{1}^{T},$
where $I$ is the identity matrix. Hence, using a bit linear algebra, we find that $\lambda_{2}(A)=\frac{k}{n(n-1)(n-k)}$. Hence, $R(S,T)=\sqrt{n\lambda_{2}(A)}=\sqrt{\frac{k}{(n-1)(n-k)}}$. ∎
Next, we give the proof of Theorem 1.7 by induction on $n$. For $n=1$, $S$ and $T$ are independent. Hence, $R(S,T)=0$ and Theorem 1.7 holds. Suppose that Theorem 1.7 holds for $n\leq N$ and consider $n=N+1$. If $m=n$ or $m=0$, then $T$ is independent of $S$. Hence, $R(S,T)=0$ and Theorem 1.7 holds. Therefore, we may assume that $0<m<n$ in the following. Let $U$ be a random subset of $[n]$ with cardinality $n-1$ containing $T$. Moreover, assume that $U$ is uniform given $T$. In this way, we have $S\subset T\subset U$ and $(S,T,U)$ is uniform. Let $\varphi$ be a measurable function such that $E\varphi(S)=0$ and $E[(\varphi(S))^{2}]=1$. Since $(S,T,U)$ is Markov, i.e. $S$ and $U$ are conditionally independent given $T$, we have
$E(\varphi(S)|T)=E(\varphi(S)|T,U).$
Write $\varphi(S)=g(U)+f(S,U)$, where
$g(U)=E(\varphi(S)|U).$
Then, we have that $E(f(S,U)|U)=0$ and $E(\varphi(S)|T,U)=g(U)+E(f(S,U)|T,U)$. Note that
$(E(\varphi(S)|T))^{2}$ $=(E(\varphi(S)|T,U))^{2}$
$=(g(U))^{2}+(E(f(S,U)|T,U))^{2}+2g(U)E(f(S,U)|T,U).$
Since $E(E(f(S,U)|T,U)|U)=E(f(S,U)|U)=0$, we have that
$E[(E(\varphi(S)|T))^{2}|U]=(g(U))^{2}+E[(E(f(S,U)|T,U))^{2}|U].$ (43)
Since $E(f(S,U)|U)=0$ and $\operatorname{Var}(f(S,U)|U)<\infty$, by (2), we have that
$E[(E(f(S,U)|T,U))^{2}|U]\leq(R(S,T|U))^{2}E[(f(S,U))^{2}|U],$ (44)
where $R(S,T|U)$ is the maximal correlation coefficient of the conditional distribution of $(S,T)$ given $U$. By the induction hypothesis, we have
$(R(S,T|U))^{2}=\frac{k(n-1-m)}{m(n-1-k)}.$ (45)
Note that
$E[(f(S,U))^{2}|U]=E[(\varphi(S)-g(U))^{2}|U]=E[(\varphi(S))^{2}|U]-(g(U))^{2}.$ (46)
By (43), (44), (45) and (46), we obtain that
$E[(E(\varphi(S)|T))^{2}|U]\leq\frac{k(n-1-m)}{m(n-1-k)}E[(\varphi(S))^{2}|U]+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)(g(U))^{2}.$
By taking the expectation on both sides, we get that
$E[(E(\varphi(S)|T))^{2}]\leq\frac{k(n-1-m)}{m(n-1-k)}E[(\varphi(S))^{2}]+\left(1-\frac{k(n-1-m)}{m(n-1-k)}\right)E[(g(U))^{2}].$ (47)
Note that $g(U)=E[\varphi(S)|U]$. Hence, by (2) and Lemma 5.2, we get that
$E[(g(U))^{2}]\leq(R(S,U))^{2}E[(\varphi(S))^{2}]=\frac{k}{(n-1)(n-k)}E[(\varphi(S))^{2}].$ (48)
Finally, combining (47) and (48), using the identity $\frac{k(n-1-m)}{m(n-1-k)}+(1-\frac{k(n-1-m)}{m(n-1-k)})\frac{k}{(n-1)(n-k)}=\frac{k(n-m)}{m(n-k)}$, we get that
$E[(E(\varphi(S)|T))^{2}]\leq\frac{k(n-m)}{m(n-k)}E[(\varphi(S))^{2}].$
Since this holds for all measurable $\varphi$ with $E\varphi(S)=0$ and $\operatorname{Var}\varphi(S)=1$, we conclude that
$(R(S,T))^{2}\leq\frac{k(n-m)}{m(n-k)}.$
To prove the converse inequality $R(S,T)\geq\sqrt{\frac{k(n-m)}{m(n-k)}}$, we take $\varphi(S)=1_{1\in S}$ and $\psi(T)=1_{1\in T}$. Then, we have that
$R(S,T)\geq\rho(\varphi(S),\psi(T))=\sqrt{\frac{k(n-m)}{m(n-k)}}.$
5.4 Proof of Corollary 1.8
We will prove Corollary 1.8 in this subsection. In the proof, we use Theorem 1.6, Theorem 1.7 and Lemma 2.1 (submultiplicative property).
Firstly, since $\partial$ is different from the possible values of $X_{i}$, we have that
$R((Y_{1},Y_{2},\ldots,Y_{n}),(Z_{1},Z_{2},\ldots,Z_{n}))=R((S,X_{S}),(T,X_{T})).$
Let $U=S\cap T$. Then, we observe that $(S,X_{S}),(U,X_{U}),(T,X_{T})$ is Markov. Hence, by Lemma 2.1, we have that
$R((S,X_{S}),(T,X_{T}))\leq R((U,X_{U}),(S,X_{S}))R((U,X_{U}),(T,X_{T})).$ (49)
We explicitly calculate $R((U,X_{U}),(S,X_{S}))$ and $R((U,X_{U}),(T,X_{T}))$ in the following lemma:
Lemma 5.3.
Recall that $(U,S)$ is uniformly distributed with the constraints that $|U|=\ell$, $|S|=k$ and $U\subset S$. We have that
$R((U,X_{U}),(S,X_{S}))=\sqrt{\frac{\ell}{k}}.$ (50)
Similarly, we have that
$R((U,X_{U}),(T,X_{T}))=\sqrt{\frac{\ell}{m}}.$ (51)
Proof.
We only give the proof of (50) as the proof of (51) is exactly the same. By Theorem 1.6, we have that
$R((U,X_{U}),(S,X_{S}))=\max(R(U,S),\max(r_{j}:j\in[n])),$
where $r_{j}$ is the best constant $r$ for the following inequality
$\sum_{u,s:j\in u\cap s}P(U=u,S=s)\alpha_{u}\beta_{s}\leq r\sqrt{\sum_{u:j\in u}P(U=u)\alpha_{u}^{2}}\sqrt{\sum_{s:j\in s}P(S=s)\beta_{s}^{2}}$ (52)
for arbitrary real constants $\alpha_{u}$ and $\beta_{s}$. By Theorem 1.7, we see that
$R(U,S)=\sqrt{\frac{\ell(n-k)}{k(n-\ell)}}\leq\sqrt{\frac{\ell}{k}}.$
Hence, it suffices to prove that
$r_{j}=\sqrt{\frac{\ell}{k}}.$ (53)
By performing the change of variables $\widetilde{\alpha}_{u}=\sqrt{P(U=u)}\alpha_{u}$ and $\widetilde{\beta}_{s}=\sqrt{P(S=s)}\beta_{s}$, we find that $r_{j}$ is exactly the spectral radius of the matrix $A=(A_{us})_{u,s:j\in u\cap s}$, where
$A_{us}=\frac{P(U=u,S=s)}{\sqrt{P(U=u)P(S=s)}}$
with the convention that $/0=0$ 0. Let $B=AA^{*}$, where $A^{*}$ is the transpose of $A$. Then, $B$ is real, symmetric and positively definite. Moreover, $r_{j}^{2}$ is exactly the maximal eigenvalue of $B$. Note that
$B_{uv}$ $=\sum_{s:j\in s}A_{us}A_{vs}$
$=\sum_{s:j\in s}\frac{P(U=u,S=s)P(U=v,S=s)}{\sqrt{P(U=u)P(U=v)}P(S=s)}$
$=\frac{\sqrt{P(U=u)}}{\sqrt{P(U=v)}}\sum_{s:j\in s}P(S=s|U=u)P(U=v|S=s).$
Then, we see that $B$ is similar to $C$, where
$C_{uv}=\sum_{s:j\in s}P(S=s|U=u)P(U=v|S=s)$
with the convention that the conditional probability is zero if it is not well-defined. Hence, $r_{j}^{2}$ is equal to the maximal eigenvalue of $C$. Note that $P(j\in U|S=s)=\ell/k$ for $j\in s$ and $P(j\in S|U=u)=1$ for $j\in u$. Hence, we have that
$\sum_{v:j\in v}C_{uv}$ $=\sum_{s:j\in s}\sum_{v:j\in v}P(S=s|U=u)P(U=v|S=s)$
$=\sum_{s:j\in s}P(S=s|U=u)P(j\in U|S=s)$
$=\frac{\ell}{k}\sum_{s:j\in s}P(S=s|U=u)$
$=\frac{\ell}{k}P(j\in S|U=u)$
$=\frac{\ell}{k}.$
Note that $C$ is a matrix with non-negative elements such that the sum of each row is the constant $\ell/k$. By Perron-Frobenius theorem, the maximal eigenvalue of $C$ is precisely $\ell/k$. Hence, we obtain that $r_{j}=\sqrt{\ell/k}$ and the proof is complete. ∎
By (49), (50) and (51), we obtain the upper bound
$R((S,X_{S}),(T,X_{T}))\leq\ell/\sqrt{km}.$
To prove the lower bound, we take a measurable function $h(x)$ such that $Eh(X_{j})=0$ and $\operatorname{Var}h(X_{j})=1$. Take $\varphi(S,X_{S})=1_{j\in S}h(X_{j})$ and $\psi(T,X_{T})=1_{j\in T}h(X_{j})$. Then, we get that
$R((S,X_{S}),(T,X_{T}))\geq\rho(\varphi(S,X_{S}),\psi(T,X_{T}))=\ell/\sqrt{km}.$
Remark 5.2.
Under the assumption of Theorem 1.6, by similar arguments as in Lemma 5.3, we have the following observation on $r_{j}$: If $P(j\in S|U=u)$ does not depend on $u$ when $j\in u$ and $P(j\in U|S=s)$ does not depend on $s$ when $j\in s$, then we have that
$r_{j}^{2}=P(j\in S|U=u)P(j\in U|S=s).$
5.5 Proof of Theorem 1.9
We prove Theorem 1.9 in this subsection.
For a non-empty subset $s\subset[n]$, we define $U_{s}=\sum_{i\in s}X_{i}$ and denote by $f_{s}(u)$ the density of $U_{s}$. Let $\rho_{s}(u)=f^{\prime}_{s}(u)/f_{s}(u)$ be the score function of $U_{s}$. Then the Fisher information
$I(U_{s})=E[\rho^{2}_{s}(U_{s})].$
For two nested subsets $s\subset t$, we have that
$E[\rho_{s}(U_{s})|U_{t}]=\rho_{t}(U_{t})$
by [MB07, Lemma 1] for the convolution identities for scores. Define
$\varphi(s,x_{s})=\lambda_{s}\rho_{s}(u_{s}),$
where $u_{s}=\sum_{i\in s}x_{i}$. Then, we have that
$E[\varphi(S,X_{S})|T=t,X_{t}]=\left(\sum_{s\subset[n]}P(S=s|T=t)\lambda_{s}\right)\rho_{t}(U_{t})=\mu_{t}\rho_{t}(U_{t}).$
Hence, we get that
$E[(E[\varphi(S,X_{S})|T=t,X_{t}])^{2}|T=t]=\mu_{t}^{2}I(U_{t}).$
Consequently, we have that
$E[(E[\varphi(S,X_{S})|T,X_{T}])^{2}]=E\left[E\left[(E[\varphi(S,X_{S})|T,X_{T}])^{2}|T\right]\right]=\sum_{t\subset[n]}P(T=t)\mu_{t}^{2}I(U_{t}).$
Note that
$E[(\varphi(S,X_{S}))^{2}]=\sum_{s\subset[n]}P(S=s)\lambda_{s}^{2}I(U_{s}).$
By (2), we have that
$E[(E[\varphi(S,X_{S})|T,X_{T}])^{2}]\leq R^{2}E[(\varphi(S,X_{S}))^{2}].$
Therefore, we obtain that
$\sum_{t\subset[n]}P(T=t)\mu_{t}^{2}I(U_{t})\leq R^{2}\sum_{s\subset[n]}P(S=s)\lambda_{s}^{2}I(U_{s}).$
5.6 Examples
Firstly, we consider an example studied in [BS25]. They have calculated the maximal correlation for the bivariate Marshall Olkin exponential distribution. We state their result in the following theorem:
Theorem 5.4 (Bücher-Staud).
Let $W_{1},W_{2}$ and $W_{3}$ be independent exponential random variables. For $i=1,2,3$, the parameter of $W_{i}$ is $\lambda_{i}>0$. Let $V_{1}=\min(W_{1},W_{3})$ and $V_{2}=\min(W_{2},W_{3})$. Then, we have that
$R(V_{1},V_{2})=\frac{\lambda_{3}}{\sqrt{(\lambda_{1}+\lambda_{3})(\lambda_{2}+\lambda_{3})}}.$
We will give an alternative proof of the upper bound
$R(V_{1},V_{2})\leq\frac{\lambda_{3}}{\sqrt{(\lambda_{1}+\lambda_{3})(\lambda_{2}+\lambda_{3})}}$ (54)
by using Corollary 1.8 and Lemma 2.3.
We take i.i.d. random variables $X_{1},X_{2},\ldots,X_{n}$. Each $X_{i}$ is an exponential random variable with parameter $\lambda$. Let $(S,T)$ be the pair of random subsets of $[n]$ defined in Corollary 1.8. Let $W_{1}=\min_{i\in S\setminus T}X_{i}$, $W_{2}=\min_{i\in T\setminus S}X_{i}$ and $W_{3}=\min_{i\in S\cap T}X_{i}$. Since $(S,T)$ is independent of $(X_{i})_{i\in[n]}$, $W_{1}$, $W_{2}$ and $W_{3}$ are independent. Moreover, $W_{1}$ is an exponential random variable of parameter $(k-\ell)\lambda$, $W_{2}$ is an exponential random variable of parameter $(m-\ell)\lambda$ and $W_{3}$ is an exponential random variable of parameter $\ell\lambda$. Then, $V_{1}=\min(W_{1},W_{3})=\min_{i\in S}X_{i}$ and $V_{2}=\min(W_{2},W_{3})=\min_{i\in T}X_{i}$. By Corollary 1.8, we have that
$R(V_{1},V_{2})\leq R((S,X_{S}),(T,X_{T}))=\frac{\ell}{\sqrt{mk}}.$
Hence, (54) holds with $\lambda_{1}=(k-\ell)\lambda$, $\lambda_{2}=(m-\ell)\lambda$ and $\lambda_{3}=\ell\lambda$. Therefore, (54) holds for rational $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. For general $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$, take $\lambda^{(N)}_{i}=[N\lambda_{i}]/N$ for $i=1,2,3$ and $N\geq 1$. For each $N\geq 1$, the corresponding random variables are $W^{(N)}_{1},W^{(N)}_{2},W^{(N)}_{3},V^{(N)}_{1}$ and $V^{(N)}_{2}$. As $N\to\infty$, $(V^{(N)}_{1},V^{(N)}_{2})$ converges in distribution to $(V_{1},V_{2})$. By Lemma 2.3, we have that
$R(V_{1},V_{2})$ $\leq\liminf_{N\to\infty}R(V^{(N)}_{1},V^{(N)}_{2})$
$\leq\liminf_{N\to\infty}\frac{\lambda_{3}^{(N)}}{\sqrt{(\lambda_{1}^{(N)}+\lambda_{3}^{(N)})(\lambda_{2}^{(N)}+\lambda_{3}^{(N)})}}$
$=\frac{\lambda_{3}}{\sqrt{(\lambda_{1}+\lambda_{3})(\lambda_{2}+\lambda_{3})}}.$
By similar arguments, we actually have the following upper bound:
Proposition 5.5.
Let $X_{1},X_{2},\ldots,X_{n}$ be i.i.d. (real) random variables. Let $1\leq\ell\leq m\leq n$. Then, we have that
$R\left(\min_{i:1\leq i\leq m}X_{i},\min_{j:\ell+1\leq j\leq n}X_{j}\right)\leq\frac{m-\ell}{\sqrt{m(n-\ell)}}.$ (55)
6 Open problems
  1. 1.
    Let $(X_{t},Y_{t})_{t\in[0,1]}$ be a two-dimensional Lévy bridge. Is there an expression for the maximal correlation coefficient $R((X_{t})_{t\geq 0},(Y_{t})_{t\geq 0})$?
  2. 2.
    By [BS25, Corollary 2.2], the upper bound in (55) is sharp if $X_{i}$ follows exponential distribution. Is it also sharp for other distributions?